A Bibliography of Publications of Henk A. van der Vorst

Henk A. van der Vorst Universiteit Utrecht

Department of Mathematics

NL-3508 TA Utrecht The Netherlands

Tel: +31 30 533732 FAX: +31 30 518394

E-mail: vorst@math.uu.nl (Internet)

19 September 2023 Version 2.46

Abstract

This bibliography records publications of Henk A. van der Vorst.

Title word cross-reference

2 [vdV86b, vdV87c]. 3 [SvdV89, vdV87a, vdV89c]. A [vdV87b, vdVM90, vdV00b]. Ax = b [vdVM90]. BiCGstab(l) [SvdVF94]. ℓ [SFvdV94]. f(A)x = b [vdV87b, vdV00b]. LU [vdV83c]. M [MvdV77]. p [BvdV01a]. R^{n} [Ano02].

-cyclic [BvdV01a]. -decompositions [vdV83c]. -Matrix [MvdV77].

1 [vdVvK83, vdV85]. 10P [vdV89b]. 1980 [Hin82]. 1989 [AK90]. 1991 [FA91, FF93, Shi91]. 1992 [HW93]. 1993 [BCEP94, SSS93]. 1998 [I^{+}99]. 1999 [F^{+}00].

2000 [vdVDE^{+}02]. 2002 [BČF04, vdVDE^{+}03]. 205 [vdVvK83, vdV85, vdV89b]. 20th [GvdV00, SvdV00a]. 23rd [JKB96].

3rd [VGH01].

4th [FA91].

5 [PMG^{+}94].

8th [BAWH00].

’98 [I^{+}99].

Abraham [vdV05]. Accelerated [FSvdV98a]. Acta [Ano93]. Advances [GGL93, GGL94, I^{+}99]. advection [VK93]. aims [van04]. Algebra [BDD^{+}95, BDDD95, BdG92, DHvdV92, DHvdV93a, DDSvdV98, DHvdV93b, HvdV00, vdVD90]. algebraic [BDDR00, VvdV87b, WS98]. Algorithm [vdVS96, CDvdV91]. Algorithms [DBGvdV02, FSvdV98b, WS98, vdVBDP01, HW93, KSV97, K^{+}04, vdVD90, BvdV96a, tDv87, vdV94a]. Alive [vdVG97]. Alternatives [HvdV03]. Amsterdam [BAWH00, tDv87]. Analysis [Ano02, DW97, vdV86a, vdV87d, Eva83, HvdV00, RvdVtM03, RvdVtM04, mMvdV02, vDHvdV00a]. Application [KvdVPG93, KvdVPG95, BMTvdV94, CDvdV91]. Applications [BvdV01a, FF93, VVSV96, Eva83, HW93, I^{+}99, Nol87, RvdVtM03, RvdVtM04, SAD^{+}00, WS98, tDv87, tRvdV95, SvdVR08]. Applied [F^{+}00]. approaches [SvdV95b, SvdV96c]. Approximate [CvdV94, CvdV97, PPvdV93, PPvdV95, BSvdV99]. April [BdG92, GH94]. Arising [vdV81, VvdV87b]. arithmetic [SvdV94b, SvdV95a]. Arnoldi [BMTvdV94, KvdVPG93, KvdVPG95]. Art [DW97, LS87, VvdVV92, vdV94b]. Aspects [tDv87, CvdV93, SvdVR08, vdV88a, vdV89e, vdV93b, vdVS98]. August [F^{+}00, I^{+}99, VGH01]. Austria [SSS93]. Automatic [VvdV77a]. Avoiding [GMS99].

Barrett [Ano95]. based [FF93]. basis [BMTvdV94]. behavior [HvdV89b, HvdV90, VvdV87a, vdV89a]. behaviour [HvdV89a, van90, vdVV93]. Belgium [BdG92]. benchmark [DEvdV01]. Benchmarks [DG93]. Berry [Ano95]. Bertin [KvdV94]. between [VvdVV92, vdV94b]. BI [vdV92a, DvdV91, SvdVF94, SvdV96a, vdV91b, vdV93a, vdV94d, vdVS98]. BI-CG [vdV92a, SvdVF94, SvdV96a, vdV93a, vdV94d, vdVS98]. BI-CGSTAB [vdV92a, DvdV91, vdV91b]. BiCG [SvdV95b, SvdV96c]. BICGSTAB [SvdV94b, SFvdV94, SvdV95a]. bidiagonal [SvdV89]. Bielefeld [Hin82]. Block [VvdVV92, SvdV89, vdV87e]. Block-ART [VvdVV92]. Blocks [BCD^{+}93, BBC^{+}94a, BBC^{+}94b]. Book [Ano95, vdV94a]. Bounds [vDHvdV01, PPvdV93, PPvdV95, vDHvdV00b, vdEFL^{+}02]. breakdown [GMS99]. Brussels [BdG92]. Building [BCD^{+}93, BBC^{+}94b, BBC^{+}94a]. Bulgaria [I^{+}99].

C [vdV94a]. calculations [VVSV96]. Carolina [BCEP94]. Catherine [FF93]. Centenary [BCEP94]. century [GvdV00, SvdV00a]. Certain [vdV81, vdV82b]. CFD [Hab95]. CG [DvdV93a, SvdVF94, SvdV96a, dvdV95, VvdV90, van90, vdV92a, vdV93a, vdV94b, vdV94d, vdV94e, vdVS98]. CG-S [van90]. CG-type [VvdV90]. CGSTAB [DvdV91, vdV91b, vdV92a]. challenges [F^{+}00]. Chan [Ano95]. Cholesky [mMvdV00]. chromodynamics [F^{+}00]. Circuit [BvdV01a, BvdV00]. close [VvdV87a]. Closer [GvdV97]. CM [PMG^{+}94]. CM-5 [PMG^{+}94]. Cocoyoc [HK^{+}02]. Coefficient [MvdV77]. College [FF93]. Colloquium [tDv87]. Communication [CvdV93, dvdV94, DvdV93a, dvdV95]. Comparative [vdVvK83, vdV85]. Comparison [vdV94b, VvdV92]. Complex [KvdVPG93, KvdVPG95, vdVM90]. Computation [VvdV77b, FF93, GvdV00, SvdV95b, SvdV96c, vdV04]. Computational [vdV02a]. computations [vLL85]. Computer [DG93, F^{+}00]. Computers [DvdV93a, DDSv91, DvdV92b, DvdV92a, DvdV93b, DDSvdV98, dvdV94, vdV86b, SvdV89, dvdV95, tDv87, vLL85, vdV86c, vdV87e, vdV89c, vdV90, vdV97]. Computing [F^{+}00, vDHvdV00b, vDHvdV01, BAWH00, CCG97, GH94, VGH01, tRvdV95, vdV89e]. Conference [AK90, BCEP94, BAWH00, FA91, GH94, HK^{+}02, SSS93, FF93, I^{+}99, vdVDE^{+}01, vdVDE^{+}02, vdVDE^{+}03].

Conjugate [AK90, K^{+}04, vdSvdV86, vdVD88, vdV92b, CvdV93, FSvdV96, vdV86c, FSvdV94, vdVS93]. construction [BMTvdV94]. control [BSvdV99]. Convergence [vdSvdV86, vdVY00, BvdV04, HvdV89a, HvdV89b, HvdV90, SvdV94b, SvdV95a, VvdV87a, vdV89a, van90, vdVS93, vdVV93]. Converging [vdV92a]. Copper [vdVDE^{+}02, vdVDE^{+}03, vdVDE^{+}01]. Cornelius [BCEP94]. Cost [dvdV94]. Cracking [BvdV96a, BvdV96b]. craftmanship [van04]. CRAY [vdVvK83, vdV85]. CRAY-1 [vdVvK83, vdV85]. critical [vdV05]. Current [JKB96]. CYBER [vdV85, vdVvK83, vdV89b]. cyclic [BvdV01a]. Czech [JKB96].

D [SvdV89, vdV86b, vdV87a, vdV87c, vdV89c]. Data [MvdV86, MvdV88]. Davidson [BvdVMT94, BFSvdV96, FSvdV98b, GSvdV02, RvdVtM03, RvdVtM04, SvdV94a, SvdV96b, SBFvdV96, SvdVM98, SvdV00b, VVSV96]. December [BCEP94]. decomposition [HW93, HK^{+}02, MvdV01, vdVS93]. Decompositions [MvdV81, vdV82c, vdV83b, vdV83c]. definite [vdV87b]. Demmel [Ano95]. Development [DvdV99a]. Developments [DvdV99b, vdV94e]. devices [FA91, SSS93]. Differences [SvdVM00, SvdVM01]. differential [Hin82]. diffusion [VK93]. dimension [vdV04]. Discretised [vdV88b]. Distributed [DvdV93a, CvdV93, dvdV95]. distributed-memory [CvdV93]. Domain [HK^{+}02, MvdV01, HK^{+}02]. Donato [Ano95]. Dongarra [Ano95].

E. [KvdV94]. ECMI [BČF04]. effect [DvdV93a, dvdV95, vdVS93]. Effects [SvdVM00, SvdVM01]. Efficient [SFvdV94, SvdVM98, vdV02b]. eigenpairs [vdV04]. Eigenproblems [vdVS96, vdVG97, SBFvdV96, SvdVM98, SvdVW99, vdV96b]. Eigenvalue [GvdV00, HvdV03, SvdV96b, SvdV00b, VvdV77a, BDDR00, BKvdV06, BvdVMT94, BvdV96a, BvdV96b, GSvdV02, PPvdV93, PPvdV95, SvdV94a, vdV02a, BFSvdV96]. Eigenvalues [VvdV77b, vDHvdV01, vDHvdV00b, VvdV87a, vdV99, vdV08]. Eighth [HW93]. Eijkhout [Ano95]. electrical [VGH01]. electromagnetics [StM05]. element [K^{+}04]. Engineering [GPS89, HK^{+}02, VGH01, WS98]. Equations [Ano02, DvdV92b, DvdV92a, DvdV93b, FSvdV98a, MvdV81, SvdV93, SFvdV94, Hin82, vdV91a, vdV96c]. error [vdEFL^{+}02]. Errors [SvdVM00, SvdVM01, van90]. ETA [vdV89b]. ETA-10P [vdV89b]. Europe [BAWH00]. evaluation [TPV92]. Exhibition [GH94]. expansion [SvdVM98]. Experiences [vdV90]. exploration [Nol87]. Extreme [VvdV77b, vDHvdV01, vDHvdV00b].

factorization [MvdV01]. Factorizations [CvdV94, CvdV97, mMvdV01, mMvdV02]. Family [mMvdV00, vdVV91, vdVV94]. Fast [vdV92a, vdVS96]. February [GGL94]. Federal [FA91]. Fifth [GPS89, SSS93]. finite [K^{+}04, SvdV94b, SvdV95a]. formation [TPV92]. formulations [CdvdV98]. FORTRAN [SvdV90, vdV86c]. fourteenth [HK^{+}02]. fourth [I^{+}99]. free [CDvdV91, CdvdV98]. Freeman [vdV94a]. Fukuoka [Shi91]. function [vdEFL^{+}02]. Further [vdVFS93].

Galerkin [vdVM90]. GAMM [HW93]. GAMM-Seminar [HW93]. Generalized [FSvdV96, VvdV77a, vdVS96, BvdVMT94, BFSvdV96, MvdV01, SvdV94a, SBFvdV96, SvdVM98, vdV82a, FSvdV94]. geophysics [Nol87]. Germany [GH94, VGH01]. Gigantic [Bus05]. global [DvdV93a, Nol87, dvdV95]. GMRES [BvdV01a, DvdV93a, HvdV89a, HvdV89b, HvdV90, VvdV92, dvdV95, vdVV91, vdVV93, vdVV94]. GMRES-like [VvdV92]. GMRES-type [BvdV01a]. GMRESR [vdVV91, vdVV94]. Gradient [AK90, CvdV93, FSvdV96, K^{+}04, vdVD88, vdV92b, FSvdV94]. Gradients [vdSvdV86, vdV86c, vdVS93]. guide [BDDR00]. Guidelines [MvdV81].

Handbook [Ano02]. handling [TPV92]. Held [AK90, FA91, FF93, GGL94, Hin82, SSS93]. Henk [Bus05]. High [BAWH00, DDSvdV98, GH94, vdV89d, vdV04]. High-Performance [DDSvdV98, GH94]. HPCN [BAWH00]. Hybrid [SvdV96a, vdVS94, SvdVF94, SvdV95b, SvdV96c, vdVFS93, vdV94e].

ICCG [vdV82d, vdV83a, vdV86b, vdV87a, vdV87c, vdV89c]. II [I^{+}99, BvdV96b, HvdV90]. III [HvdV00]. ILU [HW93]. IMA [GGL94]. IMACS [BdG92]. images [TPV92]. implementation [vdV97, vdVS98]. Implementations [BvdV96b, vdV86c]. implicit [BSvdV99, BvdV01b, mMvdV02]. implicitly [KvdVPG95, KvdVPG93]. improvement [VVSV96]. improvements [vdVFS93]. Incomplete [CvdV94, HW93, MvdV81, mMvdV00, vdV82c, vdV83b, CvdV97, MvdV01, mMvdV01, mMvdV02, vdV83c, vdVS93]. Indefinite [SvdVM00, SvdVM01]. industrial [BČF04, SAD^{+}00]. Inexact [FSvdV98a]. informatics [JKB96]. information [vdV87b]. Institute [FA91, FF93, F^{+}00]. integration [Hin82]. interdisciplinary [F^{+}00]. International [BdG92, BCEP94, BAWH00, FA91, GH94, GPS89, HK^{+}02, SSS93, Shi91, I^{+}99, VGH01]. interview [Bus05]. Introduction [vdVS94]. Invert [KvdVPG93, KvdVPG95]. Iteration [SvdV93, SvdV96b, SvdV00b, vdV83b, vdVS94, vdV00a, SvdV94a, VVSV96, vdV83c, vdV96b, vdV99]. iterations [BSvdV99]. Iterative [BCD^{+}93, BBC^{+}94a, BBC^{+}94b, BdG92, CCG97, GGL93, GvdV97, MvdV77, SvdV00a, vdV81, vdV82b, vdV91a, vdV94c, vdVC94, vdV96a, vdV97, vdVS98, vdVY00, vdV03, van04, vdV09, DEvdV01, GGL94, VvdV90, vdV87b, vdV89a, vdVFS93, vdV93b, vdV96c, vdVC97, vdV02b, vdV04].

J [KvdV94]. Jacobi

[BvdVMT94, BFSvdV96, FSvdV98b,

GSvdV02, RvdVtM03, RvdVtM04, SvdV94a, SvdV96b, SBFvdV96, SvdVM98, SvdV00b]. January [HW93]. Japan [Shi91]. Japanese [BBC^{+}94b]. John [F^{+}00]. joint [F^{+}00]. Jülich [F^{+}00]. June [AK90].

Kiel [HW93]. Krylov

[BMTvdV94, BvdV04, PPvdV93, PPvdV95, SvdVM00, SvdVM01, dvdV94, vdV87b, vdV00a, vdVY00, vdV03, vdV09].

L [vdV94a]. Lanczos

[BCEP94, CDvdV91, CdvdV98, VvdV77a, VvdV77b, vDHvdV00b, vDHvdV01, vdV82a]. Lanczos-type [CdvdV98]. Large [FSvdV98a, SvdV93, VvdV77b, vdV82b, vdV87e, vdVS96, BvdVMT94, BvdV96a, BvdV96b, Hab95, Hin82, SAD^{+}00, VvdV87b, WS98, vdV91a, vdV96c, vdV99, vdV02a, vdV03, vdV09, BKvdV06]. large-scale [BvdV96a, BvdV96b, Hab95]. lattice [F^{+}00]. least [VvdV90]. least-squares [VvdV90]. libraries [vdV88a]. library [SvdV90]. like [VvdV92]. Linear [BDD^{+}95, BDDD95, BCD^{+}93, BBC^{+}94a, BBC^{+}94b, BdG92, DHvdV92, DHvdV93a, DDSv91, DvdV92b, DvdV92a, DvdV93b, DDSvdV98, DvdV99a, GvdV97, MvdV77, MvdV81, SvdV93, SFvdV94, SvdV96b, SvdVM00, SvdV00b, SvdVM01, vdV81, vdV82b, vdV91b, vdV92a, vdVC94, vdV96a, vdVC97, vdV08, BKvdV06, BvdV00, CDvdV91, CdvdV98, DHvdV93b, DvdV99b, Hin82, SvdV00a, SvdV05, SvdV89, SvdV94a, VvdV87b, WS98, VvdV90, vdV87d, vdV87e, vdVD89, vdV90, vdVD90, vdV92b, vdV94c, vdV96c, vdV97, vdV02b, vdV03, vdV09, HvdV00]. LU [MvdV01, vdV83b]. LU-Decompositions [vdV83b].

M [KvdV94, vdV86b, vdV87c, DvdV93a, dvdV95]. machine [CvdV93]. Magnetohydrodynamics [PMG^{+}94, BMTvdV94]. Maintaining [SvdV94b, SvdV95a]. manual [SvdV90]. March [GGL94]. Massively [tRvdV95]. Mathematics [FF93, BČF04]. Matrices [BvdV01a, VvdV77b, vDHvdV01, vDHvdV00b, vdV99, vdV04]. Matrix [FSvdV98b, MvdV77, vdV81, vdV82b, vdV00b, SAD^{+}00, vdV87b]. May [BAWH00]. Medical [TPV92]. memoriam [KvdV94, vdV05]. Memory [DvdV93a, DDSv91, CvdV93, dvdV95]. Mesh [Bv91]. Method [BvdV01a, MvdV77, MvdV86, SvdV96b, SvdV00b, vDHvdV01, vdV82b, vdV86a, BMTvdV94, BvdVMT94, CvdV93, GSvdV02, GMS99, KvdVPG95, MvdV88, SvdV94a, SvdVM98, VVSV96, VvdV77b, vDHvdV00b, vdV87d, vdV87b, vdVM90, KvdVPG93]. Methods [AK90, BBC^{+}94a, BBC^{+}94b, BdG92, GGL93, GPS89, SvdV93, SvdV96a, dvdV94, vdV81, vdV83a, vdV87a, vdVS94, vdV96a, vdVY00, vdV00b, BFSvdV96, BvdV04, CCG97, CdvdV98, Eva83, GGL94, HK^{+}02, I^{+}99, K^{+}04, RvdVtM03, RvdVtM04, StM05, SvdV94b, SvdVF94, SvdV95a, SvdV95b, SBFvdV96, SvdV96c, VK93, VvdV92, vdEFL^{+}02, VvdV90, vdV82d, vdVD88, vdV89a, vdV89c, vdVV91, vdV91a, vdV92b, vdVFS93, vdV93b, vdVV94, vdV94c, vdV94e, vdV96c, vdV97, vdVC97, vdVS98, vdV99, vdV02a, vdV02b, vdV03, van04, vdV04, vdV09]. Mexico [HK^{+}02]. MHD [BFSvdV96, KvdVPG93, KvdVPG95]. MHD-eigenvalue [BFSvdV96]. Milovy [JKB96]. mind [vdV05]. Minimum [vdV93a, vdV94d, BSvdV99]. Minneapolis [GGL94]. Minnesota [GGL94]. Model [SvdVR08]. modelling [DvdV91]. Modern [vdV96c, vdV04]. Modifications [vdV93a, vdV94d]. Monitoring [VvdV77a]. Mountain [vdVDE^{+}01, vdVDE^{+}02, vdVDE^{+}03]. MRCEPA [VVSV96]. MRCI [VVSV96]. Munich [GH94].

nameplates [Bus05]. nearly [BvdV01b]. need [vdV05]. nested [vdVV91, vdVV94]. Netherlands [AK90, BAWH00, tDv87]. Network [Bv91]. networking [BAWH00, GH94]. Neumann [F^{+}00]. Newton [FSvdV98a]. Nijmegen [AK90]. NMA [I^{+}99]. Non [vdV81, vdV82b]. Non-Symmetric [vdV82b, vdV81]. Nonlinear [FSvdV98a, BKvdV06]. Nonsymmetric [SFvdV94, vdV92a, CDvdV91, CdvdV98, vdV92b, vdVFS93]. North [BCEP94]. November [JKB96, Shi91]. numerica [Ano93]. Numerical [Ano02, DHvdV92, DHvdV93a, DDSvdV98, DW97, DBGvdV02, GPS89, HvdV00, VvdV77b, VvdV87b, VK93, tDv87, vDHvdV00a, vdEFL^{+}02, vdV94a, vdVBDP01, DHvdV93b, I^{+}99, KSV97, StM05, vdVD90, F^{+}00, Hin82]. NUMVEC [SvdV90].

observations

[HvdV89a, HvdV89b, HvdV90]. obtained [vdV87b]. Occur [MvdV81]. Old [vdVG97]. operator [vdEFL^{+}02]. Optimal [SvdV93]. optimized [GSvdV02]. order [SvdVR08]. organized [FF93]. other [SvdVF94]. overlap [mMvdV02, vdEFL^{+}02]. overlapped [mMvdV01]. overview [SvdV95b, SvdV96c]. Oxford [FF93].

Paige [VvdV77b]. Paige-Style [VvdV77b]. papers [GGL94]. paradigm [MvdV01]. Parallel [Bv91, BMTvdV94, DvdV93a, DHvdV92, DHvdV93a, DHvdV93b, DvdV99a, KSV97, PMG^{+}94, dvdV94, mMvdV00, mMvdV01, vLL85, vdV86a, vdV88b, vdVD90, vdV93b, BvdV00, BvdV01b, Dv98, DvdV99b, FF93, MvdV01, dvdV95, mMvdV02, tDv87, tRvdV95, vdV87d, vdV87e, vdV89e, vdV90, vdV97, FF93, vdV94a]. Parallelizable [BvdV01a, vdVS96]. ParIC [mMvdV00]. part [BvdV96b, BvdV96a]. Partition [MvdV86, MvdV88]. PDE [vdV81, vdV88b]. PDE-Problems [vdV81]. Pencils [FSvdV98b]. Performance [DvdV92b, DvdV92a, DvdV93b, DDSvdV98, vdVvK83, vdV85, BAWH00, GH94, vdV86c, vdV89d]. Petrov [vdVM90]. Phillips [vdV94a]. pole [RvdVtM03, RvdVtM04]. pole-zero [RvdVtM03, RvdVtM04]. polynomial [SBFvdV96]. positive [vdV87b]. Pozo [Ano95]. Practical [MvdV81, vdV89e, BDDR00]. practice [JKB96]. precision [SvdV94b, SvdV95a]. Preconditioned [vdV82b, BvdVMT94, vdV86c, van90, AK90]. preconditioner [vdV94d]. Preconditioners [mMvdV00, vdV93a]. Preconditioning [Dv98, Eva83, SAD^{+}00, SvdVW99, vdV82c, RvdVtM03, RvdVtM04, vdVD88, vdV89d, vdVS93, vdV99]. Preconditionings [vdV83b, MvdV01, vdV83c]. Preface [BKvdV06, tRDvdV91a, tRDvdV91b]. presence [VvdV87a, van90]. Probabilistic [vDHvdV01, vDHvdV00b]. Problem [HvdV03]. Problems [BDD^{+}95, BDDD95, GGL94, MvdV81, SvdV96b, SvdV00b, VvdV77a, vdV81, vdV86b, vdV87a, BDDR00, BKvdV06, BvdV00, BvdVMT94, BvdV96a, BvdV96b, BFSvdV96, GSvdV02, Hab95, SAD^{+}00, SvdV94a, VvdV87b, VK93, VvdV90, vdV87c, vdV89c, vdV02a]. Proceedings [AK90, BAWH00, GPS89, SSS93, BdG92, FA91, FF93, GH94, HW93, Hin82, I^{+}99, VGH01, BCEP94, Shi91, JKB96]. processes [FA91, SSS93]. Processors [tDv87]. Progress [BČF04]. Projection [vdV00b, vdV08]. properties [SvdV94b, SvdV95a]. pseudo [mMvdV01, mMvdV02]. pseudo-overlap [mMvdV02]. pseudo-overlapped [mMvdV01]. pt [Ano02].

QCDd [vdEFL^{+}02]. QR [FSvdV98b]. Quadratic [HvdV03]. quantum [F^{+}00]. Quotient [HvdV03]. QZ [FSvdV98b].

Raleigh [BCEP94]. Rate [vdSvdV86]. Rayleigh [HvdV03]. recurrence [vdVD89]. Reducing [DvdV93a, dvdV95]. Reduction [FSvdV98b, dvdV94, SvdVR08]. Related [vdV87a, vdV89c]. Relation [VvdVV92]. relations [vdVD89]. Reliable [SvdV96a, vdV02b]. Replacement [vdVY00]. Republic [JKB96]. research [SvdVR08]. Residual [vdV93a, vdVY00, BSvdV99, vdV94d]. Residuals [SvdV96a, vdVY00]. Results [VvdV77b]. Review [vdV94a]. Reviews [Ano95]. Ritz [BvdV04, VvdV87a]. Romine [Ano95]. Rounding [SvdVM00, SvdVM01, van90].

S [van90]. scale [BKvdV06, BvdV96a, BvdV96b, Hab95, WS98]. scheme [BvdV01b, vdV82a]. Schemes [FSvdV98a, VvdV77a, BSvdV99]. Schwarz [GSvdV02]. Science [F^{+}00, HK^{+}02, WS98]. scientific [CCG97, vdV89e, VGH01]. Seismic [Nol87]. seismology [Nol87]. semiconductor [DvdV91, FA91, SSS93]. Seminar [HW93, JKB96]. Sept [SSS93]. September [FA91, FF93]. Sets [MvdV81]. Shared [DDSv91]. Shift [KvdVPG93, KvdVPG95]. shift-and-invert [KvdVPG95]. Sign [vdEFL^{+}02]. Sign-function [vdEFL^{+}02]. Simple [vdV83a]. Simulation [BvdV01a, FA91, SSS93, BvdV00, FA91, SSS93]. SIRT [VvdVV92, VvdV90, vdV94b]. SIRT- [VvdV90]. Skew [VvdV77a]. Skew-Symmetric [VvdV77a]. Sluis [vdV05]. Smoothly [vdV92a]. Sofia [I^{+}99]. SOFSEM’96 [JKB96]. software [vdV88a]. Solution [Ano02, BCD^{+}93, BBC^{+}94a, BBC^{+}94b, DvdV99a, GvdV97, Hab95, MvdV77, SvdV05, vdV81, vdV82b, vdV86a, vdV88b, vdV92a, vdV00b, BDDR00, DvdV99b, SvdV00a, VvdV87b, VvdV90, vdV87d, vdV87b, vdV89a, vdV91a, vdV96c, van04]. solutions [PPvdV93, PPvdV95]. Solver [SFvdV94, BvdV00, DEvdV01]. Solvers [GvdV97, SvdVM00, SvdVM01, vdVC94, vdVC97]. Solving [Bv91, DDSv91, DvdV92b, DvdV92a, DvdV93b, MvdV81, SvdV89, BvdVMT94, CDvdV91, vdV87b, vdVM90]. Some [HvdV89a, HvdV89b, HvdV90, vdV83a, VvdV92, vdV82d, vdV89a]. Sparse [DvdV92b, DvdV92a, DvdV93b, GGL94, VvdV77b, vdV81, vdV82b, vdVC94, SAD^{+}00, VvdV87b, VvdV90, vdV90, vdVC97]. Special [StM05]. Spectral [mMvdV02]. Spring [Hin82]. squared [CDvdV91, FSvdV96, FSvdV94]. squares [VvdV90]. St [FF93]. Stability [BSvdV99]. Stabilized [vdV83c, vdV83b]. Stable [SFvdV94, SvdV95b, SvdV96c]. Standard [DvdV92b, DvdV92a, DvdV93b, SvdVM98]. State [DW97, LS87]. state-of-the-art [LS87]. stepping [BSvdV99, BvdV01b]. Still [vdVG97]. Strategies [vdVY00]. Strategy [KvdVPG93, KvdVPG95].

Structured [GGL94]. Style

[FSvdV98b, VvdV77b]. subdomains [mMvdV01]. Subspace [vdV96b, vdV99, vdV00a, vdVY00, BMTvdV94, vdV87b]. subspaces [PPvdV93, PPvdV95, SvdVM98]. supercomputers [vdV91a]. Supercomputing [LS87, Shi91]. superlinear [vdVV93]. Surprisingly [SFvdV94]. Switzerland [FA91]. Symmetric [MvdV77, SvdVM00, SvdVM01, VvdV77a, vDHvdV01, vdV82b, vDHvdV00b, vdV81, vdV87b, vdVM90]. Symposium [BdG92, GPS89, Shi91]. System [Bv91, vdVC94, BvdV00, vdVC97]. Systems [BCD^{+}93, BBC^{+}94a, BBC^{+}94b, DDSv91, DvdV99a, FSvdV98a, MvdV77, SvdV93, SvdVM00, SvdVM01, vdV81, vdV82b, vdV86a, vdV91b, vdV92a, vdV96a, CDvdV91, CdvdV98, DvdV99b, Hin82, SvdV00a, SvdV05, SvdV89, VvdV87b, WS98, vdV87d, vdV87e, vdV90, vdV91a, vdV92b, vdV94c, vdV96c, vdV97, vdV02b, vdV03, vdV08, vdV09].

T [vdV94a]. Tchebycheff

[vdV82b, vdV83c, vdV83b]. Technical [SSS93]. Techniques [DvdV92b, DvdV92a, DvdV93b, Ano02, Hab95, SAD^{+}00]. Technology [FA91]. Templates [BDD^{+}95, BDDD95, BDDR00, BBC^{+}94a, BBC^{+}94b]. Tests [vdVvK83, vdV85]. theory [HW93, JKB96, SvdVR08]. time [BSvdV99, BvdV01b]. time-stepping [BvdV01b]. tomographic [VvdV87b]. tomography [Nol87]. tools [van04]. Topic [vdVBDP01]. Transport [MvdV86, MvdV88]. Transpose [CdvdV98, CDvdV91]. Transpose-free [CdvdV98, CDvdV91]. Transputers [Bv91]. Trends [DvdV99a, JKB96, DvdV99b]. Triangular [Bv91]. Tridiagonal [vdV86a, vdV87d, vdV87e]. True [vdVY00]. two [Hin82]. type [BvdV01a, CdvdV98, SBFvdV96, VvdV90, vdVD88, vdVM90, vdV92b, vdVS98].

University [F^{+}00, Hin82, SSS93]. Unsymmetric [vdV91b, vdV96a]. Updated [SvdV96a, KvdVPG95, KvdVPG93]. Usage [MvdV81]. Use [vdV91b]. Using [DvdV92b, DvdV92a, DvdV93b, vdV87b].

values [BvdV04, VvdV87a]. Variant [vdV92a, vdV82d]. Various [DvdV92b, DvdV92a, DvdV93b]. Vector [DDSv91, vdV86b, SvdV89, tDv87, vdV86c, vdV87e, vdV89c, vdV90]. Vectorcomputers [vdV87a, vdV87c]. Vectorial [vdV88a]. vectorizable [vdV82d]. Vectorization [vdV83a, vdVD89]. versus [vdV89b]. Very [vdVS96, vdV99]. Vienna [SSS93]. Vol. [HvdV00]. volume [FA91, StM05]. Vorst [Ano95, Bus05, GMS99].

Wang [MvdV88]. Wang’s [MvdV86]. Warnemünde [VGH01]. where [vdVM90]. Which [MvdV77]. Workshop [GGL94, F^{+}00, VGH01]. workshops [Hin82]. Wuppertal [F^{+}00].

Years [vdVG97].

zero [RvdVtM03, RvdVtM04]. Zürich [FA91].

References

Axelsson:1990:PCG

[AK90] Owe Axelsson and Lily Yu.

Kolotilina, editors. Preconditioned Conjugate Gradient Methods: Proceedings of a Conference Held in Nijmegen, the Netherlands, June 19–21, 1989, volume 1457 of Lecture Notes in Mathematics.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1990. ISBN 0-387-53515-2 (New York), 3-540-53515-2 (Berlin). LCCN QA3 .L28 no. 1457.

Anonymous:1993:AN

[Ano93] Anonymous, editor. Acta numerica 1993. Cambridge University Press, Cambridge, UK, 1993. ISBN 0-521-44356-3. ISSN 0962-4929 (print), 1474-0508 (electronic). 326 pp. LCCN QA297 1993.

Anonymous:1995:BRB

[Ano95] Anonymous. Book reviews: 26. Barrett, Berry, Chan, Demmel, Donato, Dongarra, Eijkhout, Pozo, Romine, and van der Vorst. Mathematics of Computation, 64(211):??, July 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Anonymous:2002:HNA

[Ano02] Anonymous, editor. Handbook of Numerical Analysis: Solution of Equations in R^{n} (pt. 4): techniques of …, volume VIII. North-Holland, Amsterdam, The Netherlands, 2002. ISBN 0-444-50906-2. LCCN QA297 .H287 1990; QA297 .H34 1990; QA297 .H287 1989.

Bubak:2000:HPC

[BAWH00] Marian Bubak, Hamideh Afsarmanesh, Roy Williams, and Bob Hertzberger, editors. High performance computing and networking: 8th International Conference, HPCN Europe 2000, Amsterdam, The Netherlands, May 8–10, 2000: Proceedings, volume 1823 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNCSD9. ISBN 3-540-67553-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1823. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t1823.htm;   http:
//www.springerlink.com/link.
asp?id=fundvl8nd68e;   http:
//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=1823.

Barrett:1994:TSLa

[BBC^{+}94a] Richard Barrett, Michael Berry, Tony F. Chan, James Demmel, June Donato, Jack Dongarra, Victor Eijkhout, Roldan Pozo, Charles Romine, and Henk van der Vorst. Templates for the Solution of Linear Systems: Building Blocks for Iterative Methods. SIAM, Philadelphia, PA, USA, 1994. ISBN 0-89871-328-5. xiii + 112 pp. LCCN QA297.8 .T45

1994. URL ftp://cs.utk.
edu/linalg/templates.ps.

Barrett:1994:TSLb

[BBC^{+}94b] Richard Barrett, Michael Berry, Tony F. Chan, James W. Demmel, June Donato, Jack Dongarra, Victor Eijkhout, Roldan Pozo, Charles Romine, and Henk van der Vorst. Templates for the Solution of Linear Systems: Building Blocks for Iterative Methods (Japanese). Asakura Shoten, Tokyo, Japan, 1994. ISBN 4-254-11401-X. ???? pp. LCCN ???? See book review [Ano95].

Barrett:1993:BBI

[BCD^{+}93] R. Barrett, T. Chan, J. Demmel, J. Donato, J. Dongarra, V. Eijkhout, V. Pozo, Romime C., and H. van der Vorst. Building blocks for iterative solution of linear systems. in preparation, 1993.

Brown:1994:PCL

[BCEP94] J. David Brown, Moody T. Chu, Donald C. Ellison, and Robert J. Plemmons, editors. Proceedings of the Cornelius Lanczos International Centenary Conference, Raleigh, North Carolina, December 12– 17, 1993, volume 73 of Proceedings in Applied Mathematics. SIAM, Philadelphia, PA, USA, 1994. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993.

Buikis:2004:PIM

[BČF04] Andris Buikis, Raimondas Čiegis, and Alistair Fitt, editors. Progress in industrial mathematics at ECMI 2002, volume 5 of European consortium for mathematics in industry. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-40113-X (hardcover). LCCN TA329 .E96 2002.

Bai:1995:TLAa

[BDD^{+}95] Z. Bai, D. Day, J. Demmel, J. Dongarra, M. Gu, A. Ruhe, and H. van der Vorst. Templates for linear algebra problems. LAPACK Working Note 106, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, October 1995. URL http:

//www.netlib.org/lapack/
lawns/lawn106.ps;   http:
//www.netlib.org/lapack/
lawnspdf/lawn106.pdf. UT-

CS-95-311, October 1995. Published in [BDDD95].

Bai:1995:TLAb

[BDDD95] Z. Bai, D. Day, J. Demmel, and J. Dongarra. Templates for linear algebra problems. Lecture Notes in Computer Science, 1000:115–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://www.netlib.
org/utk/papers/etemplates.
ps;   http://www.netlib.
org/utk/papers/etemplates/
paper.html.

Bai:2000:TSA

[BDDR00] Zhaojun Bai, James Demmel, Jack Dongarra, and Axel Ruhe, editors. Templates for the solution of algebraic eigenvalue problems: a practical guide. Software, Environments, and Tools. SIAM, Philadelphia, PA, USA, 2000. ISBN 0-89871-471-0 (paperback), 0-89871-958-5 (e-book). xxx + 410 pp. LCCN QA193 .T46

2000. URL http://www.loc.
gov/catdir/enhancements/
fy0726/00059507-d.html;
http://www.loc.gov/catdir/
enhancements/fy0726/00059507-

t.html. A practical guide.

Beauwens:1992:IML

[BdG92] R. Beauwens and P. de Groen, editors. Iterative Methods in Linear Algebra: proceedings of the IMACS International Symposium on Iterative Methods in Linear Algebra, Brussels, Belgium, 2–4 April, 1991. Elsevier Science Publishers, Amsterdam, The Netherlands, 1992. ISBN 0-444-89248-6. LCCN QA184 .I44 1991. Proceedings IMACS Symp. Iterative Methods in Linear Algebra, Brussels, 1991.

Booten:1996:JDM

[BFSvdV96] A. Booten, D. Fokkema, G. Sleijpen, and H. van der Vorst. Jacobi–Davidson methods for generalized MHDeigenvalue problems. Zeitschrift für Angewandte Mathematik und Mechanik, 76(suppl. 1): 131–134, 1996. CODEN ZA-MMAX. ISSN 0044-2267 (print), 1521-4001 (electronic). ICIAM/GAMM 95 (Hamburg,

Bai:2006:PLS

[BKvdV06] Zhaojun Bai, Andrew Knyazev, and Henk A. van der Vorst. Preface: Large scale linear and nonlinear eigenvalue problems. Linear Algebra and its Applications, 415(1):1–2, May 1, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Booten:1994:PAM

[BMTvdV94] J. G. L. Booten, P. M. Meijer, H. J. J. Te Riele, and H. A. van der Vorst. Parallel Arnoldi method for the construction of a Krylov subspace basis: an application in magnetohydrodynamics. In Gentzsch and Harms [GH94], pages 196–201. CODEN LNCSD9. ISBN 3-540-57981-8 (Berlin: v. 2: paperback), 0-387-57981-8 (New York: v. 2: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1– 2 (c1994). DM96.00. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t0797.htm;   http:
//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=797. Two vol-

umes.

Botchev:1999:SCA

[BSvdV99] Mikhail A. Botchev, Gerard L. G. Sleijpen, and Henk A. van der Vorst. Stability control for approximate implicit time stepping schemes with minimum residual iterations. Applied Numerical Mathematics: Transactions of IMACS, 31(3):239– 253, November 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

                                        URL http://www.elsevier.
                                        nl/gej-ng/10/10/28/46/19/
                                        17/abstract.html;   http:
                                        //www.elsevier.nl/gej-ng/
                                        10/10/28/46/19/17/article.
                                        pdf.

Bus:2005:GNI

[Bus05] Evelien Bus. Gigantic nameplates— interview with Henk van der Vorst. Nieuw Archief voor Wiskunde. Vijfde Serie, 6 (1):76–80, 2005. CODEN NAWIA7. ISSN 0028-9825.

Bisseling:1991:PTS

[Bv91] Rob H. Bisseling and Johannes G. G. van de Vorst. Parallel triangular system solving on a mesh network of transputers. SIAM J. Scient. Stat. Comput., 12(4):787–799, July 1991. CO-DEN SIJCD4. ISSN 0196-5204.

Booten:1996:CLSa

[BvdV96a] Albert Booten and Henk van der Vorst. Cracking largescale eigenvalue problems. Part I. Algorithms. Computers in Physics, 10(3):239–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://

                                        aip.scitation.org/doi/10.
                                        1063/1.168574.

Booten:1996:CLSb

[BvdV96b] Albert Booten and Henk van der Vorst. Cracking largescale eigenvalue problems, part II: Implementations. Computers in Physics, 10(4):331–??, July 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/
10.1063/1.168590.

Bomhof:2000:PLS

[BvdV00] C. W. Bomhof and H. A.

van der Vorst. A parallel linear system solver for circuit simulation problems. Numerical linear algebra with applications, 7(7–8):649–665, October/December 2000. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/73505475/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=73505475&PLACEBO=IE.pdf.

Preconditioning techniques for large sparse matrix problems in industrial applications (Minneapolis, MN, 1999).

Bomhof:2001:PGT

[BvdV01a] W. Bomhof and H. A. van der Vorst. A parallelizable GMRES-type method for pcyclic matrices, with applications in circuit simulation. In Van Rienen et al. [VGH01], pages 293–300. ISBN 3-540-42173-4 (paperback). ISSN 1439-7358. LCCN TK5 .S35 2001. Selected contributions presented at the 3rd International Workshop on Scientific Computing in Electrical Engineering.

Botchev:2001:PNI

[BvdV01b] Mike A. Botchev and Henk A. van der Vorst. A parallel nearly implicit time-stepping scheme. Journal of Computational and Applied Mathematics, 137(2):229–243, 2001. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Brandts:2004:CKM

[BvdV04] Jan Brandts and Henk van der Vorst. The convergence of Krylov methods and Ritz values. In Křížek et al. [K^{+}04], pages 47–68. ISBN 3-540-21319-8. LCCN QA218 .C66 2004.

Booten:1994:PJD

[BvdVMT94] J. G. L. Booten, H. A. van der Vorst, P. M. Meijer, and H. J. J. Te Riele. A preconditioned Jacobi–Davidson method for solving large generalized eigenvalue problems. Technical Report NM-R9414, Centrum voor Wiskunde en Informatica, P. O. Box 4079, 1009 AB Amsterdam, The Netherlands, 1994. Submitted to SIAM J. Sci. Comp.

Chan:1997:IMS

[CCG97] Raymond H. Chan, Tony F. Chan, and Gene H. (Gene Howard)

Golub, editors. Iterative methods in scientific computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 981-3083-08-5 (paperback). vii + 384 pp. LCCN QA297.8 .I835 1997.

Chan:1991:TFS

[CDvdV91] T. F. Chan, L. De Pillis, and H. A. van der Vorst. A transpose-free squared Lanczos algorithm and application to solving nonsymmetric linear systems. Technical Report Preprint 690, Utrecht University, Utrecht, The Netherlands, 1991.

Chan:1998:TFF

[CdvdV98] Tony F. Chan, Lisette de Pillis, and Henk van der Vorst. Transpose-free formulations of Lanczos-type methods for nonsymmetric linear systems. Numerical Algorithms, 17(1–2):51–66, March 1998. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.
com/content/getfile/5058/
31/12/abstract.htm;   http:
//ipsapp007.kluweronline.
com/content/getfile/5058/
31/12/fulltext.pdf.

Crone:1993:CAC

[CvdV93] L. G. C. Crone and H. A.

van der Vorst. Communication aspects of the conjugate gradient method on a distributedmemory machine. Supercomputer, 10(6):4–9, 1993. CO-DEN SPCOEL. ISSN 0168-7875.

Chan:1994:AIF

[CvdV94] Tony C. Chan and Henk A. van der Vorst. Approximate and incomplete factorizations. Technical Report 871, University of Utrecht, Department of Mathematics, 1994. To appear in: Proc of Norfolk-NASA Conference, May 1994.

Chan:1997:AIF

[CvdV97] Tony F. Chan and Henk A. van der Vorst. Approximate and incomplete factorizations. In Keyes et al. [KSV97], pages 167–202. ISBN 0-7923-4282-8. LCCN QA76.9.A43 P35 1997.

Duff:2002:NA

[DBGvdV02] I. S. Duff, W. Borchers, L. Giraud, and H. A. van der Vorst. Numerical algorithms. Lecture Notes in Computer Science, 2400:675– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/bibs/2400/24000675.
                                        htm;   http://link.springer-
                                        ny.com/link/service/series/
                                        0558/papers/2400/24000675.
                                        pdf.

Dongarra:1991:SLS

[DDSv91] Jack J. Dongarra, Iain S. Duff, Danny C. Sorensen, and Henk A. van der Vorst. Solving Linear Systems on Vector and Shared Memory Computers. SIAM, Philadelphia, PA, USA, 1991. ISBN 0-89871-270-X. x + 256 pp. LCCN QA184 .S65 1991.

Dongarra:1998:NLA

[DDSvdV98] Jack J. Dongarra, Iain S.

Duff, Danny C. Sorensen, and Henk A. van der Vorst. Numerical Linear Algebra for High-Performance Computers. Software, Environments, and Tools. SIAM, Philadelphia, PA, USA, 1998. ISBN 0-89871-428-1 (paperback), 0-89871-961-5 (e-book). xviii + 342 pp. LCCN QA76.88 .N86 1998.

Dongarra:2001:ISB

[DEvdV01] Jack Dongarra, Victor Eijkhout, and Henk van der Vorst. An iterative solver benchmark. Scientific Programming, 9(4):223–231, 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//iospress.metapress.com/
app/home/contribution.asp%
3Fwasp=64t4wprhwk589ellmv56%
26referrer=parent%26backto=
issue%2C3%2C4%3Bjournal%
2C3%2C12%3Blinkingpublicationresults%
2C1%2C1;   http://www.netlib.
org/utk/people/JackDongarra/
PAPERS/sparse-bench.pdf.

Dongarra:1993:CB

[DG93] Jack J. Dongarra and Wolfgang Gentzsch, editors. Computer Benchmarks, volume 8 of Advances in Parallel Computing. North-Holland, Amsterdam, The Netherlands, 1993. ISBN 0-444-81518-X. LCCN QA76.9.E94 C63 1993.

Demmel:1992:PNL

[DHvdV92] James W. Demmel, Michael T. Heath, and Henk A. van der Vorst. Parallel numerical linear algebra. Technical Report UCB/CSD-92-703, Computer Science Division, University of California, Berkeley, ????, 1992. ???? pp. To appear in Acta Numerica, Cambridge University Press.

Demmel:1993:PNLa

[DHvdV93a] James W. Demmel, Michael T. Heath, and Henk A. van der Vorst. Parallel numerical linear algebra. LAPACK Working Note 60, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, March 1993. URL http:

                                        //www.netlib.org/lapack/
                                        lawns/lawn60.ps;   http:
                                        //www.netlib.org/lapack/
                                        lawnspdf/lawn60.pdf. UT-

CS-93-192, March 1993. Published in [DHvdV93b].

Demmel:1993:PNLb

[DHvdV93b] James W. Demmel, Michael T. Heath, and Henk A. van der Vorst. Parallel numerical linear algebra. Acta Numerica, 2:111–197, 1993. CODEN ANUMFU. ISBN 0-521-44356-3. ISSN 0962-4929 (print), 1474-0508 (electronic).

Duff:1998:PPP

[Dv98] Iain S. Duff and Henk A.

van der Vorst. Preconditioning and parallel preconditioning. Report RAL-TR-1998-052 and TR/PA/98/23, Rutherford Appleton Laboratory and CERFACS, Chilton, Oxon, England and Toulouse, France, 1998.

Driessen:1991:BCS

[DvdV91] M. Driessen and H. van der Vorst. Bi-CGSTAB in semiconductor modelling. In Fichtner and Aemmer [FA91], pages 45–

54. ISBN 3-89191-476-8. LCCN TK7871.85 .I5767 1991.

Dongarra:1992:PVCb

[DvdV92a] Jack J. Dongarra and Henk A.

van der Vorst. Performance of various computers using standard sparse linear equations solving techniques. Supercomputer, 9(5):17–29, September 1992. CODEN SP-COEL. ISSN 0168-7875.

URL http://www.netlib.
org/utk/people/JackDongarra/
PAPERS/Performance-of-Various-
Computers-Using-Standard-
Techniques.pdf.

Dongarra:1992:PVCa

[DvdV92b] Jack J. Dongarra and Henk A.

van der Vorst. Performance of various computers using standard techniques for solving sparse linear equations. Technical Report CS-92-168, Department of Computer Science, University of Tennessee, ????, 1992. ???? pp.

DeSturler:1993:REG

[DvdV93a] Eric De Sturler and Henk A. van der Vorst. Reducing the effect of global communication in GMRES(m) and CG on parallel distributed memory computers. Technical Report 832, Department of Mathematics, University Utrecht, Utrecht, 1993.

Dongarra:1993:PVC

[DvdV93b] Jack J. Dongarra and Henk A. van der Vorst. Performance of various computers using standard sparse linear equations solving techniques. In Dongarra and Gentzsch [DG93], pages 177–188. ISBN 0-444-81518-X. LCCN QA76.9.E94 C63 1993.

deSturler:1994:CCR

[dvdV94] E. de Sturler and H. A. van der Vorst. Communication cost reduction for Krylov methods on parallel computers. In Gentzsch and Harms [GH94], pages 190–195. CODEN LNCSD9. ISBN 3-540-57981-8 (Berlin: v. 2: paperback), 0-387-57981-8 (New York: v. 2: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1– 2 (c1994). DM96.00. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t0797.htm;   http:
//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-

9743&volume=797. Two vol-

umes.

deSturler:1995:REG

[dvdV95] E. de Sturler and H. A.

van der Vorst. Reducing the effect of global communication in GMRES(m) and CG on parallel distributed memory computers. Applied Numerical Mathematics: Transactions of IMACS, 18(4):441– 459, November 9, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/apnum/cas_sub/
browse/browse.cgi?year=1995&
volume=18&issue=4&aid=627.

Duff:1999:DTPa

[DvdV99a] I. S. Duff and H. A. van der Vorst. Development and trends in the parallel solution of linear systems. Report RAL-TR-1999-027, Rutherford Appleton Laboratory, Chilton, Oxon, England, 1999. ???? pp.

Duff:1999:DTPb

[DvdV99b] Iain S. Duff and Henk A.

van der Vorst. Developments and trends in the parallel solution of linear systems. Parallel Computing, 25(13–14):1931– 1970, December 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/35/21/32/36/
39/abstract.html;   http:
//www.elsevier.nl/gej-ng/

10/35/21/32/36/39/article.
pdf.

Duff:1997:SAN

[DW97] Iain S. Duff and G. Alistair Watson, editors. The State of the Art in Numerical Analysis, volume 63 of The Institute of Mathematics and Its Applications conference series: new series. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1997. ISBN 0-19-850014-9. LCCN QA297.S775

1997. URL https://global.
oup.com/academic/product/
the-state-of-the-art-in-
numerical-analysis-9780198500148.

Based on the proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Evans:1983:PMA

[Eva83] David J. Evans, editor. Preconditioning methods: analysis and applications, volume 1 of Topics in computer mathematics. Gordon and Breach, New York, NY, USA, 1983. ISBN 0-677-16320-7. ISSN 0275-5815. xi + 556 pp. LCCN QA374 .P7 1983.

Frommer:2000:NCL

[F^{+}00] Andreas Frommer et al., editors. Numerical challenges in lattice quantum chromodynamics: joint interdisciplinary workshop of John von Neumann Institute for Computing, Jülich, and Institute of Applied Computer Science, Wuppertal University, August 1999, volume 15 of Lecture Notes in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-67732-1. ISSN 1439-7358. LCCN QC793.3.G38 N86 2000.

Fichtner:1991:SSD

[FA91] Wolfgang Fichtner and Dölf Aemmer, editors. Simulation of semiconductor devices and processes, volume 4: proceedings of the 4th International Conference on Simulation of Semiconductor Devices and Processes held at the Federal Institute of Technology, Zürich, Switzerland on September 12– 14, 1991. Hartung-Gorre Verlag, Konstanz, Switzerland, 1991. ISBN 3-89191-476-8. LCCN TK7871.85 .I5767 1991.

Fincham:1993:PCB

[FF93] A. E. Fincham and Brian J.

Ford, editors. Parallel computation: based on the proceedings of a conference on parallel computation organized by the Institute of Mathematics and Its Applications and held at St. Catherine’s College, Oxford, in September 1991, volume 46 of Conference series / The Institute of Mathematics and Its Applications. Clarendon Press, Oxford, UK, 1993. ISBN 0-19-853680-1. LCCN QA76.58 .P3755 1993.

Fokkema:1994:GCG

[FSvdV94] D. R. Fokkema, G. L. G. Sleijpen, and H. A. van der Vorst. Generalized Conjugate Gradient Squared. Technical Report 851, Department of Mathematics, University Utrecht, 1994. Submitted for Proceedings Beckenridge Conference, April, 1994.

Fokkema:1996:GCG

[FSvdV96] Diederik R. Fokkema, Gerard L. G. Sleijpen, and Henk A. van der Vorst. Generalized conjugate gradient squared. Journal of Computational and Applied Mathematics, 71(1):125– 146, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Fokkema:1998:AIN

[FSvdV98a] Diederik R. Fokkema, Gerard L. G. Sleijpen, and Henk A. van der Vorst. Accelerated inexact Newton schemes for large systems of nonlinear equations. SIAM Journal on Scientific Computing, 19(2):657– 674, March 1998. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

                                        URL http://epubs.siam.
                                        org/sam-bin/dbq/article/
                                        29614;   http://epubs.siam.
                                        org/sam-bin/dbq/toc/SISC/
                                        19/2.

Fokkema:1998:JDS

[FSvdV98b] Diederik R. Fokkema, Gerard L. G. Sleijpen, and Henk A. van der Vorst. Jacobi– Davidson style QR and QZ algorithms for the reduction of matrix pencils. SIAM Journal on Scientific Computing, 20 (1):94–125, January 1998. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/
30007;   http://epubs.siam.
org/sam-bin/dbq/toc/SISC/
20/1.

Golub:1993:RAI

[GGL93] Gene H. Golub, Anne Greenbaum, and Mitchell Luskin, editors. Recent Advances in Iterative Methods, volume 60 of The IMA volumes in mathematics and its applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-94252-1 (New York), 3-540-94252-1 (Berlin). LCCN QA297.8 .R43 1994. DM98.00. Papers from the IMA Workshop on Iterative Methods for Sparse and Structured Problems, held in Minneapolis, Minn., Feb. 24–Mar. 1, 1992.

Golub:1994:RAI

[GGL94] Gene Golub, Anne Greenbaum, and Mitchell Luskin, editors. Recent advances in iterative methods: [papers from the IMA Workshop on Iterative Methods for Sparse and Structured Problems, held in Minneapolis, Minnesota, February 24–March 1, 1992], volume 60 of The IMA Volumes in Mathematics and its Applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. ISBN 0-387-94252-1 (New York), 3-540-94252-1 (Berlin). LCCN QA297.8 .R43

                                        1994. URL http://zbmath.
                                        org/?q=an:0790.00015.

Gentzsch:1994:HPC

[GH94] Wolfgang Gentzsch and Uwe Harms, editors. High-performance computing and networking: International Conference and Exhibition, Munich, Germany, April 18–20, 1994: proceedings, volume 797 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNCSD9. ISBN 3-540-57981-8 (Berlin: v. 2: paperback), 0-387-57981-8 (New York: v. 2: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1– 2 (c1994). DM96.00. URL

                                        http://link.springer-ny.
                                        com/link/service/series/
                                        0558/tocs/t0797.htm;   http:
                                        //www.springerlink.com/openurl.
                                        asp?genre=issue&issn=0302-
                                        9743&volume=797. Two vol-

umes.

Graves-Morris:1999:ABV

[GMS99] P. R. Graves-Morris and A. Salam. Avoiding breakdown in van der Vorst’s method. Numerical Algorithms, 21(1– 4):205–223, December 1999. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic). URL http:

//ipsapp007.kluweronline.
com/content/getfile/5058/
20/7/abstract.htm;   http:
//ipsapp007.kluweronline.
com/content/getfile/5058/
20/7/fulltext.pdf.  See

[BBC^{+}94b].

Gruber:1989:PFI

[GPS89] R. (Ralf) Gruber, J. (Jacques)

Periaux, and R. P. (Richard Paul) Shaw, editors. Proceedings of the Fifth International Symposium on Numerical Methods in Engineering. Computational Mechanics, Southampton, UK, 1989. ISBN 1-85312-041-3, 0-945824-24-6, 3-540-51589-5, 0-387-51589-5. LCCN M90.E00482; QA297; M90.E01501.

Genseberger:2002:OSM

[GSvdV02] M. Genseberger, G. L. G. Sleijpen, and H. A. van der Vorst. An optimized Schwarz method in the Jacobi–Davidson method for eigenvalue problems. In Herrera et al. [HK^{+}02], pages 289–296 (electronic). ISBN 970-32-0859-2. LCCN QA402.2.

Golub:1997:CSI

[GvdV97] Gene H. Golub and Henk A.

van der Vorst. Closer to the solution: Iterative linear solvers. In Duff and Watson [DW97], pages 63–92. ISBN 0-19-850014-9. LCCN QA297.S775

1997. URL https://global.
oup.com/academic/product/

                                        the-state-of-the-art-in-
                                        numerical-analysis-9780198500148.

Based on the proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

Golub:2000:ECC

[GvdV00] Gene H. Golub and Henk A. van der Vorst. Eigenvalue computation in the 20th century. Journal of Computational and Applied Mathematics, 123(1–2):35–65, 2000. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). Numerical analysis 2000, Vol. III. Linear algebra.

Habashi:1995:STL

[Hab95] Wagdi G. Habashi, editor. Solution techniques for large-scale CFD problems, Computational methods in applied sciences. Wiley, New York, NY, USA; Chichester, UK, 1995. ISBN 0-471-95810-7. LCCN QA911 S55 1995.

Hinze:1982:NID

[Hin82] Jürgen Hinze, editor. Numerical integration of differential equations and large linear systems: proceedings of two workshops held at the University of Bielefeld, Spring 1980, volume 968 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1982. ISBN 0-387-11970-1. LCCN QA3 .L28 no. 968; QA370.

Herrera:2002:DDM

[HK^{+}02] Ismael Herrera, David E.

Keyes, et al., editors. Domain decomposition methods in science and engineering: fourteenth International Conference on Domain Decomposition Methods, Cocoyoc, Mexico. National Autonomous University of Mexico (UNAM), Mexico City, Mexico, 2002. ISBN 970-32-0859-2. LCCN QA402.2.

Huang:1989:SOCa

[HvdV89a] Y. Huang and H. A. van der Vorst. Some observations on the convergence behaviour of GMRES. Technical Report 89-09, Delft University of Technology, Faculty of Tech. Math., 1989.

Huang:1989:SOCb

[HvdV89b] Yun Qing Huang and Henk A.

van der Vorst. Some observations on the convergence behavior of GMRES. I. Natural Science Journal of Xiangtan University = Xiangtan Daxue Ziran Kexue Xuebao, 11 (4):103–116, 1989. CODEN XDZEWR. ISSN 1000-5900.

Huang:1990:SOC

[HvdV90] Yun Qing Huang and Henk A.

van der Vorst. Some observations on the convergence behavior of GMRES. II. Natural Science Journal of Xiangtan University = Xiangtan Daxue Ziran Kexue Xuebao, 12 (2):135–149, 1990. CODEN XDZEWR. ISSN 1000-5900.

Hadjidimos:2000:NAV

[HvdV00] Apostolos Hadjidimos and Henk van der Vorst, editors. Numerical analysis 2000. Vol. III. Linear algebra. Elsevier Science Publishers, Amsterdam, The Netherlands, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). i–xii and 1–531 pp. J. Comput. Appl. Math. 123 (2000), no. 1–2.

Hochstenbach:2003:ARQ

[HvdV03] Michiel E. Hochstenbach and Henk A. van der Vorst. Alternatives to the Rayleigh quotient for the quadratic eigenvalue problem. SIAM Journal on Scientific Computing, 25(2): 591–603 (electronic), March 2003. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40640.

Hackbusch:1993:IDI

[HW93] W. Hackbusch and Gabriel Wittum, editors. Incomplete decomposition (ILU): algorithms, theory, and applications: proceedings of the Eighth GAMM-Seminar, Kiel, January 24–26, 1992, volume 41 of Notes on Numerical Fluid Mechanics. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1993. ISBN 3-528-07641-0, 3-531-07641-0 (invalid checksum??). ISSN 0179-9614. LCCN QA377 .G23 1992.

Iliev:1999:RAN

[I^{+}99] O. P. (Oleg P.) Iliev et al., editors. Recent advances in numerical methods and applications II: proceedings of the fourth international conference, NMA ’98, Sofia, Bulgaria 19–23 August, 1998. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1999. ISBN 981-02-3827-4. xv + 907 pp. LCCN QA297 .R37 1999.

Jeffery:1996:STP

[JKB96] Keith (Keith G.) Jeffery, Jaroslav Král, and Miroslav Bartosek, editors. SOF-SEM’96: theory and practice of informatics: 23rd Seminar on Current Trends in Theory and Practice of Informatics, Milovy, Czech Republic, November 23–30, 1996: proceedings, volume 1175 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNCSD9. ISBN 3-540-61994-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751.S62 1996. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t1175.htm;   http:
//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=1175.

Krizek:2004:CGA

[K^{+}04] M. Křížek et al., editors. Conjugate gradient algorithms and finite element methods. Scientific computation. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. ISBN 3-540-21319-8. xv + 382 pp. LCCN QA218 .C66 2004.

Keyes:1997:PNA

[KSV97] David E. Keyes, Ahmed Sameh, and V. Venkatakrishnan, editors. Parallel numerical algorithms, volume 4 of ICASE/LaRC interdisciplinary series in science and engineering. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 0-7923-4282-8. xi + 395 pp. LCCN QA76.9.A43 P35 1997.

Kuijlaars:1994:MMJ

[KvdV94] A. Kuijlaars and H. A. van der Vorst. In memoriam: E. M. J. Bertin: 1931–1994. Nieuw Archief voor Wiskunde. Vierde Serie, 12(3):195–199, 1994. CODEN NAWIA7. ISSN 0028-9825.

Kooper:1993:AIU

[KvdVPG93] M. N. Kooper, H. A. van der Vorst, S. Poedts, and J. P. Goedbloed. Application of the Implicitly Updated Arnoldi Method with a complex shift and invert strategy in MHD. Technical Report PP 93/061, FOM, Nieuwegein, 1993.

Kooper:1995:AIU

[KvdVPG95] M. N. Kooper, H. A. van der Vorst, S. Poedts, and J. P. Goedbloed. Application of the implicitly updated Arnoldi method with a complex shift-and-invert strategy in MHD. Journal of Computational Physics, 118(2):320– 328, May 1995. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0021999185711023.

Lichnewsky:1987:SSA

[LS87] A. Lichnewsky and C. Saguez, editors. Supercomputing: stateof-the-art. North-Holland, Amsterdam, The Netherlands, 1987. ISBN 0-444-70320-9. LCCN QA76.5 .S898 1987.

mongaMade:2000:PFP

[mMvdV00] Mardochée Magolu monga Made and Henk van der Vorst. ParIC: a family of parallel incomplete Cholesky preconditioners. In Bubak et al. [BAWH00], pages 89–

98. CODEN LNCSD9. ISBN 3-540-67553-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1823. URL

http://www.springerlink.
com/link.asp?id=2lg1kttr7w3qmwfc;
http://www.springerlink.
com/openurl.asp?genre=article&
issn=0302-9743&volume=1823&
spage=89.

mongaMade:2001:PIF

[mMvdV01] Mardochée Magolu monga Made and Henk A. van der Vorst. Parallel incomplete factorizations with pseudooverlapped subdomains. Parallel Computing, 27(8):989– 1008, July ??, 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

URL http://www.elsevier.
nl/gej-ng/10/35/21/47/34/
24/abstract.html;   http:
//www.elsevier.nl/gej-ng/
10/35/21/47/34/24/article.
pdf.

mongaMade:2002:SAP

[mMvdV02] M. Magolu monga Made and Henk A. van der Vorst. Spectral analysis of parallel incomplete factorizations with implicit pseudo-overlap. Numerical linear algebra with applications, 9(1):45–64, January/February 2002. CO-DEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/88013649/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=88013649&PLACEBO=IE.pdf.

Meijerink:1977:ISM

[MvdV77] J. A. Meijerink and H. A. van der Vorst. An iterative solution method for linear systems of which the coefficient matrix is a symmetric M-matrix. Mathematics of Computation, 31(137):148–162, January 1977. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Meijerink:1981:GUI

[MvdV81] J. A. Meijerink and H. A.

van der Vorst. Guidelines for the usage of incomplete decompositions in solving sets of linear equations as they occur in practical problems. Journal of Computational Physics, 44(1):134–155, November 1981. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0021999181900413.

Michielse:1986:DTW

[MvdV86] P. Michielse and H. van der Vorst. Data transport in Wang’s partition method. Technical Report 86-32, Delft University of Technology, Delft, 1986.

Michielse:1988:DTW

[MvdV88] Peter H. Michielse and Henk A.

van der Vorst. Data transport in Wang’s partition method. Parallel Computing, 7(1):87– 95, April 1988. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Made:2001:GDD

[MvdV01] M. Magolu monga Made and H. A. van der Vorst. A generalized domain decomposition paradigm for parallel incomplete LU factorization preconditionings. Future Generation Computer Systems, 17(8):925– 932, June 2001. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.elsevier.com/
gej-ng/10/19/19/45/35/28/
abstract.html.

Nolet:1987:STA

[Nol87] Guust Nolet, editor. Seismic tomography: with applications in global seismology and exploration geophysics. Seismology and exploration geophysics. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1987. ISBN 90-277-2521-7 (paperback). x + 386 pp. LCCN QE538.5 .S429 1987.

Poedts:1994:PMC

[PMG^{+}94] S. Poedts, P. M. Meijer, J. P. Goedbloed, H. A. van der Vorst, and A. Jakoby. Parallel magnetohydrodynamics on the CM-5. In Gentzsch and Harms [GH94], pages 365–370. CODEN LNCSD9. ISBN 3-540-57981-8 (Berlin: v. 2: paperback), 0-387-57981-8 (New York: v. 2: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1– 2 (c1994). DM96.00. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t0797.htm;   http:
//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=797. Two vol-

umes.

Paige:1993:ASE

[PPvdV93] C. C. Paige, B. N. Parlett, and H. A. van der Vorst. Approximate solutions and eigenvalue bounds from Krylov subspaces. Technical Report PAM-579, Center for Pure and Appl. Math., University of California, Berkeley, CA, 1993.

Paige:1995:ASE

[PPvdV95] Chris C. Paige, Beresford N.

Parlett, and Henk A. van der Vorst. Approximate solutions and eigenvalue bounds from Krylov subspaces. Numerical linear algebra with applications, 2(2):115–133, 1995. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Rommes:2003:JDM

[RvdVtM03] J. Rommes, H. A. van der Vorst, and E. J. W. ter Maten. Jacobi–Davidson methods and preconditioning with applications in pole-zero analysis. Reports on applied and numerical analysis RANA 03-02, Department of Mathematics and Computer Science, Eindhoven University of Technology, Eindhoven, The Netherlands, 2003. 5 pp.

Rommes:2004:JDM

[RvdVtM04] J. Rommes, H. A. van der Vorst, and E. J. W. ter Maten. Jacobi–Davidson methods and preconditioning with applications in pole-zero analysis. In Buikis et al. [BČF04], pages 191–196. ISBN 3-540-40113-X (hardcover). LCCN TA329 .E96 2002.

Saad:2000:PTL

[SAD^{+}00] Yousef Saad, Owe Axelsson, Iain Duff, Wei-Pai Tang, and Andy Wathen, editors. Preconditioning techniques for large sparse matrix problems in industrial applications. Wiley, New York, NY, USA; Chichester, UK, 2000. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic). Papers from the International Conference (SPARSE ’99) held at the University of Minnesota, Minneapolis, MN, June 10–12, 1999, Numer. Linear Algebra Appl. 7 (2000), no. 7–8.

Sleijpen:1996:JDT

[SBFvdV96] Gerard L. G. Sleijpen, Albert G. L. Booten, Diederik R. Fokkema, and Henk A. van der Vorst. Jacobi–Davidson type methods for generalized eigenproblems and polynomial eigenproblems. BIT Numerical Mathematics, 36(3):595–633, September 1996. CODEN BIT-TEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (elec-

tronic). URL http://www.
mai.liu.se/BIT/contents/

bit36.html. International Linear Algebra Year (Toulouse, 1995).

Sleijpen:1994:BES

[SFvdV94] G. L. G. Sleijpen, D. R. Fokkema, and H. A. van der Vorst. BiCGstab(ℓ ): An efficient and surprisingly stable solver of nonsymmetric linear equations. In Brown et al. [BCEP94], pages 291– 293. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993.

Shimasaki:1991:PIS

[Shi91] Masaaki Shimasaki, editor.

Proceedings of the International Symposium on Supercomputing: Fukuoka, Japan, November 6–8, 1991. Kyushu University Press, Fukuoka, Kyushu, Japan, November 1991. ISBN 4-87378-284-8. LCCN QA76.88.I1991. Also published in/as Supercomputer, volume 8, number 6 (1991).

Selberherr:1993:SSD

[SSS93] Siegfried Selberherr, H. (Hannes)

Stippel, and E. (Ernst) Strasser, editors. Simulation of semiconductor devices and processes: Proceedings of the Fifth International Conference on Simulation of Semiconductor Devices and Processes held at the Technical University of Vienna, Austria, Sept. 7–9, 1993, volume 5. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 0-387-82504-5 (New York), 3-211-82504-5 (Vienna). LCCN TK7871.85 S5555 1993; TK7871.85 .S5555 1993.

Schilders:2005:SVN

[StM05] W. H. A. Schilders and E. J. W. ter Maten, editors. Special volume: numerical methods in electromagnetics, volume 13 of Handbook of numerical analysis. Elsevier, Amsterdam, The Netherlands, 2005. ISBN 0-444-51375-2. 500 pp. LCCN QA297 .H287 1989 vol. 13.

Schlichting:1989:SBB

[SvdV89] J. J. F. M. Schlichting and H. A. van der Vorst. Solving 3D block bidiagonal linear systems on vector computers. Journal of Computational and Applied Mathematics, 27(1–2):323–330, September 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0377042789903737.

Schlichting:1990:NFL

[SvdV90] J. J. F. M. Schlichting and H. A. van der Vorst. NUMVEC FORTRAN library manual. Report. Centrum voor Wiskunde en Informatica NM-9018, Stichting Mathematisch Centrum, Amsterdam, The Netherlands, October 1990. 11 pp.

Sleijpen:1993:OIM

[SvdV93] G. L. G. Sleijpen and H. A. van der Vorst. Optimal iteration methods for large linear systems of equations. In Vreugdenhil and Koren [VK93], pages 291–320. ISBN 3-528-07645-3. ISSN 0179-9614. LCCN QA901 .N86 1993. Chapter 12.

Sleijpen:1994:GJD

[SvdV94a] G. L. G. Sleijpen and H. A.

van der Vorst. A generalized Jacobi–Davidson iteration method for linear eigenvalue problems. Technical Report 856, University Utrecht, Department of Mathematics, 1994.

Sleijpen:1994:MCP

[SvdV94b] G. L. G. Sleijpen and H. A.

van der Vorst. Maintaining convergence properties of BICGSTAB methods in finite precision arithmetic. Technical Report 861, University Utrecht, Department of Mathematics, 1994.

Sleijpen:1995:MCP

[SvdV95a] Gèrard L. G. Sleijpen and Henk A. van der Vorst. Maintaining convergence properties of BiCGstab methods in finite precision arithmetic. Numerical Algorithms, 10(3–4):203– 223, October 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Sleijpen:1995:OAS

[SvdV95b] Gerard L. G. Sleijpen and Henk A. van der Vorst. An overview of approaches for the stable computation of hybrid BiCG methods. Applied Numerical Mathematics: Transactions of IMACS, 19(3):235– 254, 1995. CODEN ANMAEL.

ISSN 0168-9274 (print), 1873-5460 (electronic). Special issue on iterative methods for linear equations (Atlanta, GA, 1994).

Sleijpen:1996:RUR

[SvdV96a] G. L. G. Sleijpen and H. A. van der Vorst. Reliable updated residuals in hybrid Bi-CG methods. Computing, 56(2):141–163, 1996. CO-DEN CMPTA2. ISSN 0010-485X (print), 1436-5057 (elec-

tronic). URL http://www.
springer.at/springer.py?
Page=10&Key=362&cat=300607/
tocs/springer.py?Page=47&
Key=340&cat=3&id_abstract=
291&id_volume=26&id_journal=
8.

Sleijpen:1996:JDI

[SvdV96b] Gerard L. G. Sleijpen and Henk A. van der Vorst. A Jacobi–Davidson iteration method for linear eigenvalue problems. SIAM Journal on Matrix Analysis and Applications, 17(2):401–425, April 1996. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27042.

Sleijpen:1996:OAS

[SvdV96c] Gerard L. G. Sleijpen and Henk A. van der Vorst. An overview of approaches for the stable computation of hybrid BiCG methods. Applied Numerical Mathematics: Transactions of IMACS, 19(3):235– 254, January 15, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cgi-bin/cas/
tree/store/apnum/cas_sub/
browse/browse.cgi?year=1996&
volume=19&issue=3&aid=633.

Special issue on iterative methods for linear equations (Atlanta, GA, 1994).

Saad:2000:ISL

[SvdV00a] Yousef Saad and Henk A.

van der Vorst. Iterative solution of linear systems in the 20th century. Journal of Computational and Applied Mathematics, 123(1–2):1–33, 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). Numerical analysis 2000, Vol. III. Linear algebra.

Sleijpen:2000:JDI

[SvdV00b] Gerard L. G. Sleijpen and Henk A. van der Vorst. A Jacobi–Davidson iteration method for linear eigenvalue problems. SIAM Review, 42 (2):267–293 (electronic), June 2000. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36308.

Schenk:2005:SLS

[SvdV05] O. Schenk and H. A. van der Vorst. Solution of linear systems. In Schilders and ter Maten [StM05], pages 755–824.

ISBN 0-444-51375-2. LCCN QA297 .H287 1989 vol. 13.

Sleijpen:1994:OHB

[SvdVF94] G. L. G. Sleijpen, H. A. van der Vorst, and D. R. Fokkema. BiCGstab(l) and other hybrid Bi-CG methods. Numerical Algorithms, 7(1):75–109, June 1994. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Sleijpen:1998:EES

[SvdVM98] Gerard L. G. Sleijpen, Henk A. van der Vorst, and Ellen Meijerink. Efficient expansion of subspaces in the Jacobi– Davidson method for standard and generalized eigenproblems. Electronic Transactions on Numerical Analysis, 7:75–89 (electronic), 1998. CODEN ???? ISSN 1068-9613 (print), 1097-4067 (electronic). URL http:/

/etna.mcs.kent.edu/. Large

scale eigenvalue problems (Argonne, IL, 1997).

Sleijpen:2000:DER

[SvdVM00] Gerard L. G. Sleijpen, Henk A. van der Vorst, and Jan Modersitzki. Differences in the effects of rounding errors in Krylov solvers for symmetric indefinite linear systems. SIAM Journal on Matrix Analysis and Applications, 22(3):726– 751 (electronic), 2000. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

URL http://epubs.siam.
org/sam-bin/dbq/article/

32308;   http://epubs.siam.
org/sam-bin/dbq/toc/SIMAX/
22/3.

Sleijpen:2001:DER

[SvdVM01] Gerard L. G. Sleijpen, Henk A.

van der Vorst, and Jan Modersitzki. Differences in the effects of rounding errors in Krylov solvers for symmetric indefinite linear systems. SIAM Journal on Matrix Analysis and Applications, 22(3):726–751, July 2001. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32308.

Schilders:2008:MOR

[SvdVR08] W. H. A. Schilders, H. A.

van der Vorst, and Joost Rommes, editors. Model order reduction: theory, research aspects and applications, volume 13 of Mathematics in industry; The European Consortium for Mathematics in Industry. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 3-540-78841-7, 3-540-78840-9. LCCN TA347.L5 M63 2008.

Sleijpen:1999:PE

[SvdVW99] Gerard L. G. Sleijpen, H. A.

van der Vorst, and F. Wubs. Preconditioning for eigenproblems. In Iliev et al. [I^{+}99], pages 170–177. ISBN 981-02-3827-4. LCCN QA297 .R37 1999.

teRiele:1987:AAV

[tDv87] H. J. J. te Riele, T. J. Dekker, and H. A. van der Vorst, editors. Algorithms and applications on vector and parallel computers: Colloquium on Numerical Aspects of Vector and Parallel Processors (1985– 1986, Amsterdam, The Netherlands), volume 3(3) of Special Topics in Supercomputing. Elsevier Science Publishers, Amsterdam, The Netherlands, 1987. ISBN 0-444-70322-5 (US). LCCN QA76.6 .A45851 1987.

Todd-Pokropek:1992:MIF

[TPV92] Andrew Todd-Pokropek and M. A. Viergever, editors. Medical images: formation, handling, and evaluation, volume 98 of NATO ASI series. Series F, Computer and system sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1992. ISBN 3-540-56131-5 , 0-387-56131-5. LCCN RC78.7.D35 M42 1992; RC78.7.D35.M42; RC78.7.D35 M42 1988. Proceedings of the NATO Advanced Study Institute on the Formation, Handling, and Evaluation of Medical Images, held at Povoa de Varzim, Portugal, September 12–23, 1988.

teRiele:1991:Pa

[tRDvdV91a] Herman J. J. te Riele, T. J. Dekker, and H. A. van der Vorst. Preface. Applied Numerical Mathematics: Transactions of IMACS, 7(5):367–368, June 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
016892749190007M.

teRiele:1991:Pb

[tRDvdV91b] Herman J. J. te Riele, T. J.

Dekker, and H. A. van der Vorst. Preface. Applied Numerical Mathematics: Transactions of IMACS, 8(2):91– 92, September 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.sciencedirect.com/
science/article/pii/016892749190044Z.

teRiele:1995:MPC

[tRvdV95] H. J. J. te Riele and H. A.

van der Vorst, editors. Massively parallel computing and applications. Elsevier Science Publishers, Amsterdam, The Netherlands, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Papers from the symposia held in Amsterdam, 1993–1994, Appl. Numer. Math. 19 (1995), no. 1–2.

vanderVorst:1990:CBP

[van90] H. A. van der Vorst. The convergence behaviour of preconditioned CG and CG-S in the presence of rounding errors. Lecture Notes in Mathematics, 1457:126–136, 1990. CODEN LNMAA2. ISBN 3-540-53515-2 (print), 3-540-46746-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

                                        tronic). URL http://link.
                                        springer.com/chapter/10.
                                        1007/BFb0090905/.

vanderVorst:2004:ISM

[van04] Henk A. van der Vorst. Iterative solution methods: aims, tools, craftmanship. In Applied mathematics entering the 21st century, pages 371–392. SIAM, Philadelphia, PA, USA, 2004.

vandenEshof:2002:NMQ

[vdEFL^{+}02] J. van den Eshof, A. Frommer, Th. Lippert, K. Schilling, and H. A. van der Vorst. Numerical methods for the QCDd overlap operator. I. Sign-function and error bounds. Computer Physics Communications, 146(2):203–224, July 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0010465502004551.

vanDooren:2000:NAV

[vDHvdV00a] P. M. van Dooren, A. Hadjidimos, and H. A. van der Vorst, editors. Numerical analysis 2000, volume 3. Elsevier Science Publishers, Amsterdam, The Netherlands, 2000. ISBN 0-444-50598-9. xii + 531 pp. LCCN QA184 .L67

                                        2000. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0614/2002282450-d.html.

Linear algebra—linear systems and eigenvalues, Reprint of J. Comput. Appl. Math. 123 (2000), no. 1–2 [ MR1798517 (2001f:65003)].

vanDorsselaer:2000:CPB

[vDHvdV00b] Jos L. M. van Dorsselaer, Michiel E. Hochstenbach, and Henk A. van der Vorst. Computing probabilistic bounds for extreme eigenvalues of symmetric matrices with the Lanczos method. SIAM Journal on Matrix Analysis and Applications, 22(3):837–852, 2000. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36685.

vanDorsselaer:2001:CPB

[vDHvdV01] Jos L. M. van Dorsselaer, Michiel E. Hochstenbach, and Henk A. van der Vorst. Computing probabilistic bounds for extreme eigenvalues of symmetric matrices with the Lanczos method. SIAM Journal on Matrix Analysis and Applications, 22(3):837–852, July 2001. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36685.

vanderSluis:1986:RCC

[vdSvdV86] A. van der Sluis and H. A.

van der Vorst. The rate of convergence of conjugate gradients. Num. Math, 48(5): 543–560, May 1986. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

vanderVorst:1981:ISM

[vdV81] Henk A. van der Vorst. Iterative solution methods for certain sparse linear systems with a non-symmetric matrix arising from PDEproblems. Journal of Computational Physics, 44(1):1–19, November 1981. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0021999181900346.

vanderVorst:1982:GLS

[vdV82a] H. A. van der Vorst. A generalized Lanczos scheme. Mathematics of Computation, 39 (160):559–561, October 1982. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

vanderVorst:1982:PTI

[vdV82b] H. A. van der Vorst. A preconditioned Tchebycheff iterative solution method for certain large sparse linear systems with a non-symmetric matrix. In Hinze [Hin82], pages 323–333. ISBN 0-387-11970-1. LCCN QA3 .L28 no. 968; QA370.

vanderVorst:1982:PID

[vdV82c] H. A. van der Vorst. Preconditioning by Incomplete Decompositions. PhD thesis, Utrecht University, 1982.

vanderVorst:1982:VVS

[vdV82d] Henk A. van der Vorst. A vectorizable variant of some ICCG methods. SIAM J. Scient. Stat. Comput., 3(3):350– 356, September 1982. CODEN SIJCD4. ISSN 0196-5204.

vanderVorst:1983:VSS

[vdV83a] H. van der Vorst. On the vectorization of some simple ICCG methods. In First International Conference on Vector and Parallel Computation in Scientific Applications, Paris, page ?? ????, ????, 1983.

vanderVorst:1983:SIL

[vdV83b] H. A. van der Vorst. Stabilized incomplete LU-decompositions as preconditionings for the Tchebycheff iteration. In Evans [Eva83], page ?? ISBN 0-677-16320-7. ISSN 0275-5815. LCCN QA374 .P7 1983.

vanderVorst:1983:SID

[vdV83c] Henk A. van der Vorst.

Stabilized incomplete LUdecompositions as preconditionings for the Tchebycheff iteration. In Evans [Eva83], pages 243–263. ISBN 0-677-16320-7. ISSN 0275-5815. LCCN QA374 .P7 1983.

vanderVorst:1985:CPT

[vdV85] H. A. van der Vorst. Comparative performance tests on the CRAY-1 and CYBER 205. In van Leeuwen and Lenstra [vLL85], pages 33–54. ISBN 90-6196-297-8. LCCN QA76.6 .P3485 1985.

vanderVorst:1986:APS

[vdV86a] H. van der Vorst. Analysis of a parallel solution method for tridiagonal systems. Technical Report 86-06, Department of Mathematics and Information, Delft University of Technology, Delft, The Netherlands, 1986.

vanderVorst:1986:MIP

[vdV86b] H. van der Vorst. (M)ICCG for 2D problems on vector computers. Technical Report 86-55, Department of Mathematics and Information, Delft University of Technology, Delft, The Netherlands, 1986.

vanderVorst:1986:PFI

[vdV86c] H. A. van der Vorst. The performance of FORTRAN implementations for preconditioned conjugate gradients on vector computers. Parallel Computing, 3(1):49–58, March 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

vanderVorst:1987:IRM

[vdV87a] H. van der Vorst. ICCG and related methods for 3D problems on vectorcomputers. Technical Report A-18, Data Processing Center, Kyoto University, Kyoto, Japan, 1987.

vanderVorst:1987:ISM

[vdV87b] H. A. van der Vorst. An iterative solution method for solving f(A)x = b, using Krylov subspace information obtained for the symmetric positive definite matrix A. Journal of Computational and Applied Mathematics, 18(2):249–263, May 1987. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042787900203.

vanderVorst:1987:MIP

[vdV87c] H. A. van der Vorst. (M)ICCG for 2D problems on vectorcomputers. In Lichnewsky and Saguez [LS87], page ?? ISBN 0-444-70320-9. LCCN QA76.5 .S898 1987.

vanderVorst:1987:APS

[vdV87d] Henk A. van der Vorst. Analysis of a parallel solution method for tridiagonal linear systems. Parallel Computing, 5(3):303– 311, November 1987. CO-DEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

vanderVorst:1987:LTB

[vdV87e] Henk A. van der Vorst. Large tridiagonal and block tridiagonal linear systems on vector and parallel computers. Parallel Computing, 5(1–2):45–54, July 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Proceedings of the international conference on vector and parallel computing—issues in applied research and development (Loen, 1986).

vanderVorst:1988:VAS

[vdV88a] H. van der Vorst. Vectorial aspects of software libraries. Supercomputer, 5(1):33–41, January 1988. CODEN SPCOEL. ISSN 0168-7875.

vanderVorst:1988:PSD

[vdV88b] H. A. van der Vorst. Parallel solution of discretised PDE’s. IEEE Trans. on Magnetics, 24 (1):286–290, 1988. CODEN IEMGAQ. ISSN 0018-9464 (print), 1941-0069 (electronic).

vanderVorst:1989:CBS

[vdV89a] H. A. van der Vorst. The convergence behavior of some iterative solution methods. In Gruber et al. [GPS89], pages 61–72. ISBN 1-85312-041-3, 0-945824-24-6, 3-540-51589-5, 0-387-51589-5. LCCN M90.E00482; QA297; M90.E01501.

vanderVorst:1989:EVC

[vdV89b] H. A. van der Vorst. ETA-10P versus Cyber 205. Supercomputer, 6(1):17–22, January 1989. CODEN SPCOEL. ISSN 0168-7875.

vanderVorst:1989:IRM

[vdV89c] H. A. van der Vorst. ICCG and related methods for 3D problems on vector computers. Computer Physics Communications, 53(1–3):223–235, May 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
0010465589901628.

vanderVorst:1989:HPP

[vdV89d] Henk A. van der Vorst. High performance preconditioning. SIAM J. Scient. Stat. Comput., 10(6):1174–1185, November 1989. CODEN SIJCD4. ISSN 0196-5204. Sparse matrix algorithms on supercomputers.

vanderVorst:1989:PAP

[vdV89e] Henk A. van der Vorst. Practical aspects of parallel scientific computing. Future Generation Computer Systems, 4(4):285– 291, March 1989. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

vanderVorst:1990:EPV

[vdV90] H. A. van der Vorst. Experiences with parallel vector computers for sparse linear systems. Supercomputer, 7(3):28– 35, May 1990. CODEN SP-COEL. ISSN 0168-7875.

vanderVorst:1991:IMS

[vdV91a] H. A. van der Vorst. Iterative methods for the solution of large systems of equations on supercomputers. Adv. Water Resources, 13:137–146, 1991. CODEN AWREDI. ISSN 0309-1708 (print), 1872-9657 (electronic).

vanderVorst:1991:UBC

[vdV91b] H. A. van der Vorst. The use of Bi-CGSTAB for unsymmetric linear systems. In Shimasaki [Shi91], pages 206–212 (or 206–211??). ISBN 4-87378-284-8. LCCN QA76.88.I1991. Also published in/as Supercomputer, volume 8, number 6 (1991).

vanderVorst:1992:BCF

[vdV92a] H. A. van der Vorst. BI-CGSTAB: a fast and smoothly converging variant of BI-CG for the solution of nonsymmetric linear systems. SIAM J. Scient. Stat. Comput., 13(2):631– 644, March 1992. CODEN SI-JCD4. ISSN 0196-5204.

vanderVorst:1992:CGT

[vdV92b] H. A. van der Vorst. Conjugate gradient type methods for nonsymmetric linear systems. In Beauwens and de Groen [BdG92], pages 67–76. ISBN 0-444-89248-6. LCCN QA184 .I44 1991. Proceedings IMACS Symp. Iterative Methods in Linear Algebra, Brussels, 1991.

vanderVorst:1993:MRM

[vdV93a] Henk A. van der Vorst. Minimum residual modifications to Bi-CG and preconditioners. In Golub et al. [GGL93], pages 217–225. ISBN 0-387-94252-1 (New York), 3-540-94252-1 (Berlin). LCCN QA297.8 .R43 1994. DM98.00. Papers from the IMA Workshop on Iterative Methods for Sparse and Structured Problems, held in Minneapolis, Minn., Feb. 24–Mar. 1, 1992.

vanderVorst:1993:PAI

[vdV93b] Henk A. van der Vorst. Parallel aspects of iterative methods. In Fincham and Ford [FF93], pages 175–186. ISBN 0-19-853680-1. LCCN QA76.58 .P3755 1993.

vanderVorst:1994:BRB

[vdV94a] Henk van der Vorst. Book review: Parallel Numerical Algorithms (T. L. Freeman and C. Phillips). SIAM Review, 36(4): 678–679, December 1994. CO-DEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://www.
siam.org/journals/sirev/
sirev364.htm.

vanderVorst:1994:CBA

[vdV94b] Henk A. van der Vorst. A comparison between ART, CG and SIRT. In Brown et al. [BCEP94], pages 302–304. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993.

vanderVorst:1994:IML

[vdV94c] Henk A. van der Vorst. Iterative methods for linear systems. In Brown et al. [BCEP94], page 277. ISBN 0-89871-339-0. LCCN QC19.2 .C67 1993.

vanderVorst:1994:MRM

[vdV94d] Henk A. van der Vorst. Minimum residual modifications to Bi-CG and to the preconditioner. In Golub et al. [GGL93], pages 217–225. ISBN 0-387-94252-1 (New York), 3-540-94252-1 (Berlin). LCCN QA297.8 .R43 1994. DM98.00. Papers from the IMA Workshop on Iterative Methods for Sparse and Structured Problems, held in Minneapolis, Minn., Feb. 24–Mar. 1, 1992.

vanderVorst:1994:RDH

[vdV94e] Henk A. van der Vorst. Recent developments in hybrid CG methods. In Gentzsch and Harms [GH94], pages 174–183. CODEN LNCSD9. ISBN 3-540-57981-8 (Berlin: v. 2: paperback), 0-387-57981-8 (New York: v. 2: paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .I57 1994 v.1– 2 (c1994). DM96.00. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t0797.htm;   http:
//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=797. Two vol-

umes.

vanderVorst:1996:IMU

[vdV96a] H. A. van der Vorst. Iterative methods for unsymmetric linear systems. In Jeffery et al. [JKB96], pages 217–234. CODEN LNCSD9. ISBN 3-540-61994-1 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.751.S62 1996. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t1175.htm;   http:
//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=1175.

vanderVorst:1996:SIE

[vdV96b] Henk van der Vorst. Subspace iteration for eigenproblems. CWI Quarterly, 9(1-2): 151–160, 1996. CODEN ???? ISSN 0922-5366. SMC 50 jubilee (Amsterdam, 1996).

vanderVorst:1996:MMI

[vdV96c] Henk A. van der Vorst. Modern methods for the iterative solution of large systems of linear equations. Nieuw Archief voor Wiskunde. Vierde Serie, 14(1):127–143, 1996. CODEN NAWIA7. ISSN 0028-9825. 31st Dutch Mathematical Conference (Groningen, 1995).

vanderVorst:1997:IML

[vdV97] Henk A. van der Vorst. Iterative methods for linear systems and implementation on parallel computers. In Chan et al. [CCG97], pages 1–44. ISBN 981-3083-08-5 (paperback). LCCN QA297.8 .I835 1997.

vanderVorst:1999:SIM

[vdV99] Henk van der Vorst. Subspace iteration methods with preconditioning for eigenvalues of very large matrices. In Iliev et al. [I^{+}99], pages 124–134. ISBN 981-02-3827-4. LCCN QA297 .R37 1999.

vanderVorst:2000:KSI

[vdV00a] Henk A. van der Vorst. Krylov subspace iteration. Computing in Science and Engineering, 2(1):32–37, January/February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

URL http://dlib.computer.
org/cs/books/cs2000/pdf/
c1032.pdf;   http://www.
computer.org/cse/cs1999/
c1032abs.htm.

vanderVorst:2000:SPM

[vdV00b] Henk A. van der Vorst. Solution of f(A)x = b with projection methods for the matrix A. In Frommer et al. [F^{+}00], pages 18–28. ISBN 3-540-67732-1. ISSN 1439-7358. LCCN QC793.3.G38 N86 2000.

URL http://link.springer.
com/chapter/10.1007/978-
3-642-58333-9_2/.

vanderVorst:2002:CML

[vdV02a] Henk A. van der Vorst. Computational methods for large eigenvalue problems. In Anonymous [Ano02], pages 3–179. ISBN 0-444-50906-2. LCCN QA297 .H287 1990; QA297 .H34 1990; QA297 .H287 1989.

vanderVorst:2002:ERI

[vdV02b] Henk A. van der Vorst. Efficient and reliable iterative methods for linear systems. Journal of Computational and Applied Mathematics, 149(1):251–265, 2002. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). Scientific and engineering computations for the 21st century—methodologies and applications (Shizuoka, 2001).

vanderVorst:2003:IKM

[vdV03] Henk A. van der Vorst. Iterative Krylov methods for large linear systems, volume 13 of Cambridge Monographs on Applied and Computational Mathematics. Cambridge University Press, Cambridge, UK, 2003. ISBN 0-521-81828-1 (print), 0-511-61511-6 (ebook). xiv + 221 pp. LCCN QA297.8 .V67 2003. URL

http://ebooks.cambridge.
org/chapter.jsf?bid=CBO9780511615115.

vanderVorst:2004:MMI

[vdV04] Henk A. van der Vorst. Modern methods for the iterative computation of eigenpairs of matrices of high dimension. Zeitschrift für Angewandte Mathematik und Mechanik, 84 (7):444–451, 2004. CODEN ZAMMAX. ISSN 0044-2267 (print), 1521-4001 (electronic).

vanderVorst:2005:NCM

[vdV05] Henk van der Vorst. The need for a critical mind—in memoriam Abraham van der Sluis (1928–2004). Nieuw Archief voor Wiskunde. Vijfde Serie, 6(1):17–19, 2005. CODEN NAWIA7. ISSN 0028-9825.

vanderVorst:2008:LSE

[vdV08] Henk van der Vorst. Linear systems, eigenvalues, and projection. In Schilders et al. [SvdVR08], pages 33–45. ISBN 3-540-78841-7, 3-540-78840-9. LCCN TA347.L5 M63 2008.

vanderVorst:2009:IKM

[vdV09] Henk A. van der Vorst. Iterative Krylov methods for large linear systems, volume 13 of Cambridge Monographs on Applied and Computational Mathematics. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-18370-7. xiv + 221 pp. Reprint of the 2003 original.

vanderVorst:2001:TNA

[vdVBDP01] Henk A. van der Vorst, Rob Bisseling, Iain S. Duff, and Bernard J. Philippe. Topic 11 numerical algorithms. Lecture Notes in Computer Science, 2150:566– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.
com/link/service/series/
0558/bibs/2150/21500566.
htm;   http://link.springer-
ny.com/link/service/series/
0558/papers/2150/21500566.
pdf.

vanderVorst:1994:LSS

[vdVC94] Henk A. van der Vorst and Tony C. Chan. Linear system solvers: Sparse iterative solvers. Technical Report 869, University of Utrecht, Department of Mathematics, 1994. To appear in: Proc of Norfolk-NASA Conference, May 1994.

vanderVorst:1997:LSS

[vdVC97] Henk A. van der Vorst and Tony F. Chan. Linear system solvers: sparse iterative methods. In Chan et al. [CCG97], pages 91–118. ISBN 981-3083-08-5 (paperback). LCCN QA297.8 .I835 1997.

vanderVorst:1988:CGT

[vdVD88] Henk A. van der Vorst and Kees Dekker. Conjugate gradient type methods and preconditioning. Journal of Computational and Applied Mathematics, 24(1–2):73–87, November 1988. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0377042788903445.

vanderVorst:1989:VLR

[vdVD89] H. A. van der Vorst and K. Dekker. Vectorization of linear recurrence relations. SIAM J. Scient. Stat. Comput., 10(1): 27–35, January 1989. CODEN SIJCD4. ISSN 0196-5204.

vanderVorst:1990:PAN

[vdVD90] Henk A. van der Vorst and Paul Van Dooren, editors. Parallel algorithms for numerical linear algebra, volume 1 of Advances in Parallel Computing. North-Holland, Amsterdam, The Netherlands, 1990. ISBN 0-444-88621-4. x + 330 pp. LCCN QA76.5 .P31458 1990.

vanderVorst:2001:CMC

[vdVDE^{+}01] Henk van der Vorst, Iain Duff, Howard Elman, Roland Freund, Tim Kelley, Seymour Parter, et al., editors. Copper Mountain Conference. SIAM, Philadelphia, PA, USA, 2001. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97414. Papers

from the 6th Conference on Iterative Methods held in Copper Mountain, CO, April 3–7, 2000, SIAM J. Sci. Comput. 23 (2001), no. 2.

vanderVorst:2002:CMC

[vdVDE^{+}02] Henk van der Vorst, Iain Duff, Howard Elman, Ronal Freund, Tim Kelley, Seymour Parter, Gerhard Starke, Nick Trefethen, Panayot Vassilevski, Homer Walker, and Olof Widlund. 2000 Copper Mountain Conference. SIAM Journal on Scientific Computing, 23(2):vii, March 2002. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97414.

vanderVorst:2003:CMC

[vdVDE^{+}03] Henk van der Vorst, Iain Duff, Howard Elman, Ronal Freund, Tim Kelley, Seymour Parter, Gerhard Starke, Nick Trefethen, Panayot Vassilevski, Homer Walker, and Olof Widlund. 2002 Copper Mountain Conference. SIAM Journal on Scientific Computing, 25(2):vii, March 2003. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/97415. Held in

Copper Mountain, CO, March 24–29, 2002.

vanderVorst:1993:FIN

[vdVFS93] Henk A. van der Vorst, Diederik R. Fokkema, and Gerard L. G. Sleijpen. Further improvements in nonsymmetric hybrid iterative methods. In Selberherr et al. [SSS93], pages 89–92. ISBN 0-387-82504-5 (New York), 3-211-82504-5 (Vienna). LCCN TK7871.85 S5555 1993; TK7871.85 .S5555 1993.

vanderVorst:1997:YOS

[vdVG97] Henk A. van der Vorst and Gene H. Golub. 150 years old and still alive: Eigenproblems. In Duff and Watson [DW97], pages 93–120. ISBN 0-19-850014-9. LCCN QA297.S775

1997. URL https://global.
oup.com/academic/product/
the-state-of-the-art-in-
numerical-analysis-9780198500148.

Based on the proceedings of a conference on the state of the art in numerical analysis. Organized by the Institute of Mathematics and Its Applications and held at York University in April 1996.

vanderVorst:1990:PGT

[vdVM90] H. A. van der Vorst and J. B. M. Melissen. A Petrov– Galerkin type method for solving Ax = b, where A is symmetric complex. IEEE Trans. on Magnetics, 26(2):706–708, 1990. CODEN IEMGAQ. ISSN 0018-9464 (print), 1941-0069 (electronic).

vanderVorst:1993:EID

[vdVS93] Henk A. van der Vorst and Gerard G. L. Sleijpen. The effect of incomplete decomposition preconditioning on the convergence of Conjugate Gradients. In Hackbusch and Wittum [HW93], pages 179– 187. ISBN 3-528-07641-0, 3-531-07641-0 (invalid checksum??). ISSN 0179-9614. LCCN QA377 .G23 1992.

vanderVorst:1994:IHI

[vdVS94] H. A. van der Vorst and G. L. G. Sleijpen. An introduction to hybrid iteration methods. In W. G. Habashi, editor, Proceedings of the International Workshop on Solution Techniques for Large-Scale CFD Problems, pages 143–159. CERCA, Montréal, Québec, Canada, 1994.

vanderVorst:1996:PFA

[vdVS96] H. A. van der Vorst and G. L. G. Sleijpen. A parallelizable and fast algorithm for very large generalized eigenproblems. Lecture Notes in Computer Science, 1184:686– ??, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanderVorst:1998:IBC

[vdVS98] H. A. van der Vorst and G. L. G. Sleijpen. Iterative Bi-CG type methods and implementation aspects. In Winter Althaus and Spedicato [WS98], pages 217–253. ISBN 0-7923-4975-X (hardcover). LCCN QA214 .A44 1998.

vanderVorst:1991:GFN

[vdVV91] H. A. van der Vorst and C. Vuik. GMRESR: a family of nested GMRES methods. Technical Report 91-80, Delft University of Technology, Faculty of Tech. Math., 1991.

vanderVorst:1993:SCB

[vdVV93] H. A. van der Vorst and C. Vuik. The superlinear convergence behaviour of GM-RES. Journal of Computational and Applied Mathematics, 48(3):327–341, November 30, 1993. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
037704279390028A.

vanderVorst:1994:GFN

[vdVV94] H. A. van der Vorst and C. Vuik. GMRESR: a family of nested GMRES methods. Numerical linear algebra with applications, 1(4):369– 386, 1994. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

vanderVorst:1983:CPT

[vdVvK83] H. van der Vorst and J. van Kats. Comparative performance tests on the CRAY-1 and Cyber 205. Preprint. Published in [vdV85]., May 1983.

vanderVorst:2000:RRS

[vdVY00] Henk A. van der Vorst and Qiang Ye. Residual replacement strategies for Krylov subspace iterative methods for the convergence of true residuals. SIAM Journal on Scientific Computing, 22(3):835–852, May 2000. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35386.

VanRienen:2001:SCE

[VGH01] Ursula Van Rienen, Michael Gunther, and Dirk Hecht, editors. Scientific computing in electrical engineering: proceedings of the 3rd international workshop, August 20–23, 2000, Warnemünde, Germany, volume 18 of Lecture Notes in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. ISBN 3-540-42173-4 (paperback). ISSN 1439-7358. LCCN TK5 .S35 2001. Selected contributions presented at the 3rd International Workshop on Scientific Computing in Electrical Engineering.

Vreugdenhil:1993:NMA

[VK93] Cornelis Boudewijn Vreugdenhil and Barry Koren, editors. Numerical methods for advection–diffusion problems, volume 45 of Notes on numerical fluid mechanics. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1993. ISBN 3-528-07645-3. ISSN 0179-9614. xii + 373 pp. LCCN QA901 .N86 1993.

vanLeeuwen:1985:PCC

[vLL85] J. van Leeuwen and J. K.

Lenstra, editors. Parallel computers and computations, volume 9 of CWI syllabus. Centrum voor Wiskunde en Informatica, P. O. Box 4079, 1009 AB Amsterdam, The Netherlands, 1985. ISBN 90-6196-297-8. 184 pp. LCCN QA76.6 .P3485 1985.

VanKats:1977:AML

[VvdV77a] J. M. Van Kats and H. A.

van der Vorst. Automatic monitoring of Lanczos schemes for symmetric or skew-symmetric generalized eigenvalue problems. Technical Report TR-7, Academisch Computer Centrum Utrecht, 1977.

VanKats:1977:NRP

[VvdV77b] J. M. Van Kats and H. A.

van der Vorst. Numerical results of the Paige-style Lanczos method for the computation of extreme eigenvalues of large sparse matrices. Technical Report TR-3, Academisch Computer Centrum Utrecht, 1977.

vanderSluis:1987:CBR

[VvdV87a] A. Van der Sluis and H. A.

van der Vorst. The convergence behavior of Ritz values in the presence of close eigenvalues. Linear Algebra and its Applications, 88/89: 651–694, 1987. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

VanderSluis:1987:NSL

[VvdV87b] A. Van der Sluis and H. A. van der Vorst. Numerical solution of large sparse linear algebraic systems arising from tomographic problems. In Nolet [Nol87], chapter 3, pages 49–83. ISBN 90-277-2521-7 (paperback). LCCN QE538.5 .S429 1987.

vanderSluis:1990:SCT

[VvdV90] A. Van der Sluis and H. A. van der Vorst. SIRT- and CG-type methods for the iterative solution of sparse linear least-squares problems. Linear Algebra and its Applications, 130:257–303 (or 257– 302??), 1990. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic). Linear algebra in image reconstruction from projections.

Vuik:1992:CSG

[VvdV92] C. Vuik and H. A. van der Vorst. A comparison of some GMRES-like methods. Linear Algebra and its Applications, 160:131–162, 1992. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

VanDijke:1992:RBA

[VvdVV92] M. C. A. Van Dijke, H. A.

van der Vorst, and M. A. Viergever. On the relation between ART, block-ART and SIRT. In Todd-Pokropek and Viergever [TPV92], pages ix + 700. ISBN 3-540-56131-5 , 0-387-56131-5. LCCN RC78.7.D35 M42 1992; RC78.7.D35.M42; RC78.7.D35 M42 1988. Proceedings of the NATO Advanced Study Institute on the Formation, Handling, and Evaluation of Medical Images, held at Povoa de Varzim, Portugal, September 12–23, 1988.

VanDam:1996:IDI

[VVSV96] H. J. J. Van Dam, J. H. Van Lenthe, G. L. G. Sleijpen, and H. A. Van Der Vorst. An improvement of Davidson’s iteration method: Applications to MRCI and MRCEPA calculations. Journal of Computational Chemistry, 17(3):267– 272, February 1996. CODEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

WinterAlthaus:1998:ALS

[WS98] G. Winter Althaus and E. Spedicato, editors. Algorithms for large scale linear algebraic systems: applications in science and engineering, volume 508 of NATO ASI series. Series C, Mathematical and physical sciences. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1998.

ISBN 0-7923-4975-X (hardcover). LCCN QA214 .A44 1998.